### v1

# Towards Long-Term User Welfare in Recommender Systems via Creator-Oriented Information Revelation

Abstract.
Improving the long-term user welfare (e.g., sustained user engagement) has become a central objective of recommender systems (RS). In real-world platforms, the creation behaviors of content creators plays a crucial role in shaping long-term welfare beyond short-term recommendation accuracy, making the effective steering of creator behavior essential to foster a healthier RS ecosystem. Existing works typically rely on re-ranking algorithms that heuristically adjust item exposure to steer creators’ behavior. However, when embedded within recommendation pipelines, such a strategy often conflicts with the short-term objective of improving recommendation accuracy, leading to performance degradation and suboptimal long-term welfare.
The well-established economics studies offer us valuable insights for an alternative approach without relying on recommendation algorithmic design: revealing information from an information-rich party (sender) to a less-informed party (receiver) can effectively change the receiver’s beliefs and steer their behavior. Inspired by this idea, we propose an information-revealing framework, named Long-term Welfare Optimization via Information Revelation (LoRe). In this framework, we utilize a classical information revelation method (i.e., Bayesian persuasion) to map the stakeholders in RS, treating the platform as the sender and creators as the receivers. To address the challenge posed by the unrealistic assumption of traditional economic methods, we formulate the process of information revelation as a Markov Decision Process (MDP) and propose a learning algorithm trained and inferred in environments with boundedly rational creators. Extensive experiments on two real-world RS datasets and across different creator modeling approaches demonstrate that our method can effectively outperform existing fair re-ranking methods and information revealing strategies in improving long-term user welfare.
Recommender System, Information Asymmetry, Information Revelation, Bounded Rationality
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
1. Introduction
With the rapid growth of user-generated content (UGC) on the Web (Krumm et al., 2008), modern recommender platforms (e.g., TikTok, YouTube) have evolved into complex ecosystems where users and content creators continuously interact and influence each other (Bobadilla et al., 2013; Ye et al., 2025). In such dynamic environments, platforms are now shifting their focus from solely optimizing short-term objectives, such as maximizing users’ immediate engagement behaviors (e.g., clicks or purchases), to fostering long-term user welfare (i.e., sustained user engagement), aiming to ensure the lasting health and sustainability of the ecosystem (Xu et al., 2023; Guo et al., 2023).
To enhance the long-term user welfare of recommender systems (RS), previous works have shown that the creators’ creation behavior (i.e., creating new items) is a key factor (Ye et al., 2025; Xu et al., 2023; Yao et al., 2024) beyond recommendation accuracy (i.e., prediction accuracy of user behavior). This is because creators continuously upload new items based on previous user feedback, which shapes the item pool of RS and indirectly influences what users are exposed to in the long run (Ben-Porat and Tennenholtz, 2018; Ben-Porat et al., 2019; Boutilier et al., 2024). Considering the importance of creators’ behavior, most existing works utilize heuristic re-ranking algorithms to steer creators’ content creation. For example, they increase item exposure for small creators to enhance creator retention and stimulate content creation (Guo et al., 2023; Xu et al., 2023; Ye et al., 2024). However, previous works have shown that such a re-ranking strategy, when embedded within recommendation pipelines, often conflicts with the short-term objectives of improving accuracy, inevitably trading off recommendation performance and harming users’ short-term experience (Ye et al., 2024; Xu et al., 2023). Therefore, how to steer creators’ content creation behavior beyond relying on recommendation algorithmic design to enhance the long-term welfare remains a significant challenge in RS.
The investigation of behavioral steering among strategic agents has a long history in economics (Milgrom and Roberts, 1987; Dierkens, 1991; Mishra et al., 1998; Crawford et al., 2018), offering valuable insights for enhancing the long-term welfare of RS. In economics studies, a common approach is to leverage information revelation under conditions of information asymmetry (i.e., agents possess unequal amounts of information) to influence agents’ behavior. Such a strategy aims to design an information revelation strategy, which specifies what information the information-rich party (sender) should disclose to the less-informed party (receiver). By revealing certain information unavailable to the receiver, the receiver can update their beliefs and adjust their strategy accordingly, leading to observable behavioral changes.
Inspired by this idea, we can steer creators’ behavior to enhance long-term welfare by revealing information, viewing the platform as the sender and creators as the receivers. Such a strategy is feasible and practical in real-world platforms, because information asymmetry naturally exists between the platform and creators in RS (Prasad et al., 2023; Boutilier et al., 2024; Ye et al., 2025). Similar but simpler idea have been adopted in industrial practice, offering real-world scenarios and potential applications for more advanced information revelation approaches. For example, platforms like TikTok provide creators with suggestions to produce trending content (see Appendix A).
However, directly applying information revelation to RS is challenging. Traditional methods in economics studies (Yao et al., 2024; Prasad et al., 2023; Boutilier et al., 2024) typically assume that receivers are completely rational, i.e., they can always select the optimal strategy to maximize their utility given the available information. In contrast, real-world creators are bounded rational (Simon, 1990; Kahneman and Tversky, 2013), because their decision-making is constrained by the limited cognitive capacity, attention, and information-processing ability of humans.
To address this challenge, we propose our information-revealing framework tailored for RS, named Long-term Welfare Optimization via Information Revelation (LoRe). Specifically, we map the key components of classical information revelation methods in economics (i.e., Bayesian persuasion (Gan et al., 2022a; Kamenica and Gentzkow, 2011)) onto the corresponding stakeholders in RS. Based on this mapping, we formulate the platform’s information revelation process as a Markov Decision Process (MDP) (Puterman, 1990) and propose a learning algorithm to optimize the information revelation strategy. Instead of relying on the assumption of completely rational agents, LoRe is trained and inferred within a realistic reinforcement learning (RL) environment, where bounded rational users and creators interact with each other.
Unlike previous works that rely on heuristic exposure allocation in the re-ranking stage to steer creators’ behavior (Guo et al., 2023), LoRe serves as an independent module for the platform and does not depend on any specific recommendation algorithm. Experiments on two real-world recommendation datasets demonstrate that LoRe outperforms both creator-fair re-ranking methods and simple information-revealing strategies in improving long-term user welfare (i.e., sustained user clicks). Moreover, since LoRe operates independently, it can be seamlessly integrated with existing ranking or re-ranking algorithms to further enhance long-term welfare.
We summarize our contributions as follows:
(1) We highlight that steering creators’ behavior through information revealing provides an effective approach to enhancing long-term user welfare beyond relying on recommendation algorithmic design.
(2) Inspired by economic studies, we propose a novel information revelation method, LoRe, which models the revelation process as a Bayesian persuasion problem and learns to optimize the revelation strategy in environments where boundedly rational agents interact.
(3) Extensive experiments on real-world datasets demonstrate that our method can effectively improve the long-term user welfare of RS by steering creators’ content creation.
2. Related Work
Multi-Stakeholder RS. Multi-stakeholder RS have received significant attention in recent years in both academic research and industrial practice (Abdollahpouri et al., 2020; Chen et al., 2025). Unlike traditional RS, which primarily focuses on providing the most suitable items based on each user’s request (He et al., 2017; Rendle et al., 2012), multi-stakeholder RS also takes into account the interests of platforms and creators, aligning with the long-term goals of many multi-stakeholder real-world platforms (e.g., TikTok, YouTube). Most existing studies on multi-stakeholder RS can be broadly categorized into two directions: (1) designing better RS algorithms that aim to achieve multi-stakeholder objectives (e.g., provider fairness (Ye et al., 2024; Xu et al., 2025), platform profit (Pei et al., 2019; Azaria et al., 2013)) while maintaining recommendation accuracy; (2) incorporating creators and platforms into the RS environment to enable more comprehensive evaluation of multi-stakeholder RS. For example, some studies focused on developing re-ranking algorithms that allocate exposure resources fairly among different providers to achieve different provider fairness objectives (e.g., Gini coefficient (Do and Usunier, 2022), Max-Min Fairness (Ye et al., 2024; Xu et al., 2023)). Ye et al. (2025) introduces an LLM-based creator simulator, CreAgent, aiming to achieve a more realistic evaluation of multi-stakeholder RS. However, existing works mainly focus on developing and evaluating RS algorithms, and do not consider another crucial side, i.e., steering creators via creation suggestions to indirectly optimize the long-term welfare of RS.
Information asymmetry in RS. A relevant topic to our work is the literature considering the information asymmetry in RS (Prasad et al., 2023; Yao et al., 2023, 2024; Boutilier et al., 2024; Ye et al., 2025). Boutilier et al. (2024); Ye et al. (2025) highlight the significant information asymmetry in RS, but they do not proposed a method to improve the welfare of RS under such asymmetry. Prasad et al. (2023) propose a prompting policy that promises to provide creators a certain level of user clicks to influence the their content creation. Yao et al. (2023, 2024) suggest influencing content creation by adjusting reward weights to encourage different types of content. However, these studies model creators as completely rational agents, ignoring the bounded rationality observed in real-world creators, which limits the applicability and performance of their methods.
Information revelation under information asymmetry. Our work is also related to the literature on information revelation under information asymmetry in economics, where a prominent approach is Bayesian persuasion (Kamenica and Gentzkow, 2011). In Bayesian persuasion, an informed party (sender) designs an information revelation strategy to reveal the optimal information to a less informed party (receiver) (Kamenica and Gentzkow, 2011). The most relevant Bayesian persuasion models to our work consider one sender (the platform) and multiple receivers (the creators) (Wang, 2013; Arieli and Babichenko, 2019; Castiglioni et al., 2021; Bernasconi et al., 2023; Zhao et al., 2025), where the sender simultaneously reveals information to multiple receivers. These studies investigate how to design information disclosure strategies to maximize the sender’s utility in a static setting. There is also a stream of research on dynamic Bayesian persuasion (Gan et al., 2022b, a; Wu et al., 2022), which examines repeated interactions between a sender and a receiver. However, because of their strong assumption of completely rational receivers, these models cannot be directly applied to real-world RS scenarios, where creators typically exhibit bounded rationality.
3. Preliminaries
In this section, we first introduce the settings of real-world RS. Then we introduce the basic idea of Bayesian Persuasion. Finally, we describe how to model creators as bounded rational agents.
3.1. Recommender Systems
Stakeholders in RS. Considering a real-world content recommendation platform (e.g., YouTube, TikTok). The platform needs to recommend items (e.g., videos, articles, etc.) produced by content creators to users. Let $\mathcal{U}=\{u_{i}\}_{i=1}^{|\mathcal{U}|}$, $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$, $\mathcal{G}=\{g_{i}\}_{i=1}^{|\mathcal{G}|}$ represent the set of users, creators, and item genres that creators can produce. Besides, let $\mathcal{I}$ denote the current corpus of items that can be recommended to users.
Interactions between different stakeholders. We model the interaction between the platform, creators, and users as a dynamic process with total $T$ interaction rounds. At the beginning of each round $t\in\{1,2,\dots,T\}$, each creator $c_{j}$ will produce an item with the genre in $\mathcal{G}$. The items produced by creators in each round will be added to $\mathcal{I}$. Then, some users will arrive and access the platform. For each accessing user, the platform needs to select $k$ items from $\mathcal{I}$ for recommendation. Each user can choose to click or skip the items recommended to him. After that, the platform records the interaction between users and items. Specifically, we use $\bm{Y}$ to represent the click matrix, where $\bm{Y}_{i,j,t}=1$ if item $i\in\mathcal{I}$ is clicked by user $u_{j}$ in round $t$ and 0 otherwise. In addition, we use $\mathcal{H}_{c_{j}}=\{(i,k_{i})|i\in\mathcal{I}\}$ to denote historical information of items created by creator $c_{j}$, where $(i,k_{i})\in\mathcal{H}_{c_{j}}$ means that the item $i\in\mathcal{I}$ produced by creator $c_{j}$ has been clicked $k_{i}$ times. For easy notation, let $\mathcal{H}=\{\mathcal{H}_{c_{1}},\mathcal{H}_{c_{2}},\dots,\mathcal{H}_{c_{|\mathcal{C}|}}\}$ denote the set of all creators’ historical information.
Information asymmetry in RS. In real-world RS, the platform has the historical information $\mathcal{H}$ of all creators and the interaction information $\bm{Y}$ of all users and items. In contrast, each creator $c_{j}$ only has its own historical information $\mathcal{H}_{c_{j}}$. Obviously, there is a significant information asymmetry between platform and creators. Therefore, how to improve the long-term welfare under information asymmetry is an unexplored and urgent problem to be addressed.
Figure 1. The overall framework of our method LoRe, which consists two phases: Online Inference and Offline Training.
3.2. Bayesian Persuasion
Information asymmetry formulation. The standard Bayesian persuasion considers an idealized model with a sender and a receiver. Both the receiver’s utility function $\mu_{r}(a_{r},w)$ and the sender’s utility function $\mu_{s}(a_{r},w)$ depend on receiver’s action $a_{r}\in\mathcal{A}_{r}$ and state $w\in\mathcal{W}$. The state $w$ follows a cumulative distribution function $F(w)$ with probability density function $f(w)$. It is assumed that both sender and receiver know $F(w)$, but the sender exactly knows the real state $w$ after $w$ is sampled. Thus, the sender have more information than the receiver, which causes the information asymmetry.
Bayesian persuasion under information asymmetry. The receiver will choose an optimal action $a_{r}^{*}$ to maximize her expected utility, and the sender will also obtain a corresponding utility. However, the sender can use a so-called signaling scheme to reveal the receiver some information, and then influence receiver’s action. Specifically, we use $\sigma=(\mathcal{A}_{s},\pi)$ to denote a signaling scheme, where $\mathcal{A}_{s}$ is the set of possible signals (messages) and $\pi:\mathcal{W}\times\mathcal{A}_{s}\mapsto[0,1]$ is the signaling strategy which maps the state to a distribution of signals. For instance, $\pi(a_{s}|w)$ denotes the probability of sending signal $a_{s}\in\mathcal{A}_{s}$ when the real state is $w$. According to the revelation principle (Myerson, 1979), each signal $a_{s}$ can be viewed as an action suggestion for receiver, that is $|\mathcal{A}_{r}|=|\mathcal{A}_{s}|$, where $a_{s}\in\mathcal{A}_{s}$ means suggesting receiver taking corresponding action $a_{r}\in\mathcal{A}_{r}$ and receiver will follows sender’s suggestion [^1] Some constraints can be used to ensure that the receiver will follow the sender’s suggestion, and these constraints do not affect the optimal solution. For details, please refer to the reference (Kamenica and Gentzkow, 2011).. By designing the optimal strategy $\pi^{*}$, the sender can improve both the utility of sender and receiver under information asymmetry (Kamenica and Gentzkow, 2011).
Table 1. Correspondence between elements in the Bayesian Persuasion and RS. See Section 4 for undefined notations.
Bayesian Persuasion Recommender System
sender platform
receivers creators
state $w$ content distribution $\bm{e}$ and trust $\bm{d}$
signal set $\mathcal{A}_{s}$ suggestion set $\mathcal{A}_{p}$
signaling strategy $\pi$ information revealing strategy $\pi$
3.3. Bounded Rationality Modeling
Motivated by the advantage of Bayesian persuasion under information asymmetry, we adopt the idea of this framework to guide how the platform reveals information in RS. Simply put, we allow the platform to send suggestions to creators, i.e., suggesting creators creating specific genres of content, which is called the platform’s information revelation strategy.
In the standard Bayesian persuasion, receivers are assumed to be completely rational, which means they can precisely choose an action $a_{r}^{*}$ to maximize their utility $\mu_{r}(a_{r},w)$. This also allows senders to design signaling strategy that ensure that receivers completely follow the sender’s action suggestion. However, creators in reality are typically bounded rational, as they cannot compute a specific utility function to maximize their utility. Thus, platform in RS cannot make sure that creators will follow platform’s suggestion. To model creators’ bounded rationality, inspired by Prasad et al. (2023), we design a trust-based model, where each creator $c_{i}$ has a trust degree $d_{{}_{i}}\in[0,1]$. When creator $c_{i}$ receives platform’s suggestion, the probability that creator $c_{i}$ follows the suggestion is $d_{{}_{i}}$ .
If creator $c_{i}$ dose not follow platform’s suggestion, we assume that he can use any strategy that does not perform utility calculation to choose an action, such as randomly selecting from the genres of previously created content, or selecting the most popular genre of previously created content, etc. That is, we do not restrict the creators’ decision model, and any decision model can be embedded in our approach that is introduced in next section. In the experiment section, to verify the universality of our method for different creator decision models, we use different model to simulate creators’ decision-making, such as LLM-based simulation model and heuristic-based simulation models.
We correspond the elements of RS with those in the Bayesian persuasion, as shown in Table 1.
$\bullet$ The platform and creators are regarded as the sender and receivers, respectively. This is because the platform, who possesses more information, can guide the creator’s creation by revealing information.
$\bullet$ The content distribution in RS and creators’ trust degree are viewed as the state in Bayesian persuasion.
$\bullet$ The suggestion that platform sends to the creators is considered as the signal in Bayesian persuasion.
$\bullet$ The platform’s strategy of sending suggestions (information revelation strategy $\pi$) is corresponds to the signaling strategy $\pi$ in Bayesian persuasion.
4. Our Approach: LoRe
In this section, to take into account creators’ bounded rationality and optimize platform’s information revelation strategy, we propose a novel method, LoRe. The overall algorithm workflow of LoRe is shown in Figure 1 and Algorithm 1. To optimize the long-term welfare of RS, we first formulate the dynamic interaction process under information asymmetry between the platform and creators as a MDP, which is described in Section 4.1. Then, we design a RL algorithm to optimize the platform’s information revelation strategy under such a dynamic environment, as presented in Section 4.2.
4.1. MDP for Platform and Creators Interactions
In our method, we use $\mathcal{A}_{p}=\{a_{i}\}_{i=0}^{|\mathcal{G}|}$ to denote the set of platform’s suggestions, where $a_{0}$ means no suggestion and $a_{i}\ (i\neq 0)$ means that platform suggests the creator to produce item with genre $g_{i}$. At the beginning of each round $t$, the platform will send suggestions selected from $\mathcal{A}_{p}$ to creators. Note that different creators may receive different suggestions. After receiving the suggestion, each creator will decide whether to follow the suggestion and choose to produce an item. Then, for each arriving user, the platform will generate $k$ items to recommend and each user can choose to click or skip these items. Then, the MDP for the interaction between platform and creators can be defined as a tuple $M=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where:
$\bullet$ State $\mathcal{S}$: $\mathcal{S}$ is the set of states. Each state is denoted by
$\bm{s}^{t}=(e_{1}^{t},e_{2}^{t},\dots,e_{|\mathcal{C}|}^{t};\hat{d}_{1}^{t},\hat{d}_{2}^{t},\dots,\hat{d}_{|\mathcal{C}|}^{t})\in\mathcal{S},$
which contains the distribution of item genres and the predicted creators’ trust in round t1t-1, where $e_{i}^{t}$ denotes the genre of item created by creator $c_{i}$ and $\hat{d}_{i}^{t}$ represents the predicted trust of creator $i$ in round t1t-1. We use $\bm{e}=(e_{1},e_{2},\dots,e_{|\mathcal{C}|})$ to denote content distribution and use $\bm{\hat{d}}=(\hat{d}_{1},\hat{d}_{2},\dots,\hat{d}_{|\mathcal{C}|})$ to represent the predicted trust vector. In simple terms, we have $\bm{s}^{t}=(\bm{e}^{t};\bm{d}^{t})$.
$\bullet$ Action $\mathcal{A}$: $\mathcal{A}$ is the set of the platform’s actions. Platform’s each action $\bm{a}^{t}\in\mathcal{A}$ is a vector $(a_{1}^{t},a_{2}^{t},\dots,a_{|\mathcal{C}|}^{t})$, where $a_{i}^{t}\in\mathcal{A}_{p}$ represents the suggestion sent to creator $c_{i}$ in round $t$.
$\bullet$ Transition function $\mathcal{T}$: The transition probability function is defined as $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$, where $\mathcal{T}(\bm{s}^{t+1}|\bm{s}^{t},\bm{a}^{t})$ denotes the probability that the state at round $t+1$ is $\bm{s}^{t+1}$ given that the the state at round $t$ is $\bm{s}^{t}$ and the platform takes action $\bm{a}^{t}$.
$\bullet$ Reward function $\mathcal{R}$: The reward function is defined as $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$, where $\mathcal{R}(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1})$ denote the total user clicks (the measure of welfare) in round $t$, given that state at round $t$ is $\bm{s}^{t}$, the platform takes action $\bm{a}^{t}$, and the state transitions to $\bm{s}^{t+1}$ in round $t+1$.
$\bullet$ Discount rate $\gamma$: $\gamma\in[0,1]$ is the discount factor that quantifies the present value of future rewards.
Note that we incorporate predicted creators’ trust degree into the state, which can help capture creators’ bounded rational behavior. The platform’s strategy of sending suggestion signals is defined as $\pi:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$, where $\pi(\bm{a}^{t}|\bm{s}^{t})$ is the probability of taking action $\bm{a}^{t}$ when the state at round $t$ is $\bm{s}^{t}$. Then, the expected discounted user clicks under strategy $\pi$ can be formulated as:
(1) $J(\pi)=\operatorname{\mathbb{E}}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1})\right],$
where $\tau$ is a trajectory, i.e., $\tau=(\bm{s}^{0},\bm{a}^{0},\bm{s}^{1},\bm{a}^{1},\dots)$ and $\tau\sim\pi$ is a shorthand indicating that the distribution over trajectories depends on $\pi:\bm{a}^{t}\sim\pi(\cdot|\bm{s}^{t})$, $\bm{s}^{t+1}\sim\mathcal{T}(\cdot|\bm{s}^{t},\bm{a}^{t})$. Furthermore, we respectively use $V^{\pi}(\bm{s}^{t})$, $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ and $A^{\pi}(\bm{s}^{t},\bm{a}^{t})$ to denote the state value function, state-action value function and the advantage function, where $V^{\pi}(\bm{s}^{t})$ is the value of state $\bm{s}^{t}$, $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ is the value of taking action $\bm{a}^{t}$ under state $\bm{s}^{t}$ and $A^{\pi}(\bm{s}^{t},\bm{a}^{t})=Q^{\pi}(\bm{s}^{t},\bm{a}^{t})-V^{\pi}(\bm{a}^{t})$ [^2] In practice, the advantage function $A^{\pi}(\bm{s}^{t},\bm{a}^{t})$ is calculated by the generalized advantage estimation (Schulman et al., 2015).. $V^{\pi}(s^{t})$ and $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ are common functions in MDP, and their definitions can be found in appendix B.
4.2. Reinforcement Learning Algorithm for Platform’s Strategy
In this section, we aim to optimize the long-term welfare in Eq(1) under the dynamic user-creator-platform interaction environment by learning an information-revealing strategy $\pi$. To achieve this, we design an actor-critic based RL algorithm. The overall workflow of our algorithm is shown in Algorithm 1, which contains two phases: the online inference (Line 4-13) and the offline training (Line 14-19).
0: Discount rate $\gamma$, clip hyperparameter $\epsilon$, learning rate lrlr, total rounds $T$, training epochs $M$, cumulative data rounds NN, reply buffer $\mathcal{B}=\{(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})\}$.
0: The policy loss $L(\pi_{\theta})$ decreases by less than 1% over 10 consecutive rounds.
1: Initialize state $\bm{s}^{1}$ from log data.
2:for $t=1$ to $T/N$ do
3:  Clear reply buffer: $\mathcal{B}=\{\}$.
4:  // $\texttt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} Phase 1: Online Inference}}$.
5:  for $r=1$ to NN do
6:   Get item genres’ distribution $\bm{e}^{t}$ and predict trust $\hat{\bm{d}}^{t}$.
7:   Get state $\bm{s}^{t}=(\bm{e}^{t},\hat{\bm{d}}^{t})$.
8:   Sample action $\bm{a}^{t}\sim\pi_{\theta}(\bm{s}^{t})$. $\texttt{Platform sends suggestions}$.
9:   $\texttt{Creators produce items}$.
10:   The next state transitions to $\bm{s}^{t+1}$.
11:   $\texttt{Users interact with items (click or skip)}$.
12:   Get the total users’ clicks $\mathcal{R}$.
13:   Add $(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ to reply buffer $\mathcal{B}$.
14:  end for
15:  // $\texttt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} Phase 2: Offline Training}}$.
16:  for $i=1$ to $M$ do
17:   Calculate actor loss $L$ and critic loss $L^{TD}$ based on $\mathcal{B}$ according to Eq. (2) and Eq. (3).
18:   Update actor network using the gradient $\nabla_{\pi_{\theta}}L$.
19:   Update critic network using the gradient $\nabla_{V_{\phi}}L^{TD}$.
20:  end for
21:end for
Algorithm 1 The overall workflow of our method LoRe.
4.2.1. Online Inference
At the beginning, the platform gets the current item genres’ distribution and predicts creators’ trust degree, which together constitute the state $\bm{s}^{t}$. We show how to predict creators’ trust degree in the following subsection. Then, the actor network $\pi_{\theta}$, which is parameterized by $\theta$, takes $\bm{s}^{t}$ as input and outputs the action distribution $\bm{P}^{t}=\pi_{\theta}(\bm{s}^{t})$, where the input $\bm{s}^{t}$ and output $\bm{P}^{t}$ are illustrated as follows.
Input. As shown in Figure 1, we represent the state $\bm{s}^{t}\in\mathbb{R}^{|\mathcal{C}|\times(|\mathcal{G}|+2)}$ as a one-hot matrix, where each row of the matrix is the one-hot vector of the item created by corresponding creator connected with the corresponding trust degree. The length of the one-hot vector is $|\mathcal{G}|+2$, because there are cases where the creator dose not create item.
Output. In the classic actor network, the actor usually outputs a distribution over the action space $\mathcal{A}$. However, the number of possible actions is $|\mathcal{A}|=(|\mathcal{G}|+1)^{|\mathcal{C}|}$, which is too large to handle. To address this problem, we design the actor to output the distribution over suggestion set $\mathcal{A}_{p}$ rather than the distribution over action set $\mathcal{A}$. That is, instead of outputting a joint distribution over all platform’s possible actions, we output an independent distribution over the suggested actions for each creator. Specifically, the output of actor is $\bm{P}=\pi_{\theta}(\bm{s}^{t})\in\mathbb{R}^{|\mathcal{C}|\times(\mathcal{G}|+1)}$, where each row of $\bm{P}$ is a distribution over the suggestion set for corresponding creator. Then, the suggestions sent to creators are sampled from the corresponding row of $\bm{P}$, which constitute the platform’s action $\bm{a}^{t}\in\mathbb{R}^{|\mathcal{C}|}$ as shown in Figure 1.
After sending suggestions to creators, creators produce items and the genres of these new items with creators’ trust degree constitute the next state $\bm{s}^{t+1}$. After that, platform recommends items to each user and get the total users’ clicks $\mathcal{R}$ as reward. Finally, the data $b^{t}=(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ is then added to the reply buffer $\mathcal{B}$ which is used to train the actor and critic networks in the offline training phase.
4.2.2. Offline Training
We use the collected reply buffer $\mathcal{B}$ to train the actor and critic networks. We denote the critic network by $V_{\phi}(\bm{s})$, which is parameterized by $\phi$. The critic network takes state $\bm{s}^{t}$ as input and outputs the estimated state value of $\bm{s}^{t}$, which is used to approximate the state value function $V^{\pi}(\bm{s})$ and help optimize the actor network.
Actor update. To update the actor, we use the clipped surrogate objective (Schulman et al., 2017) as the loss function:
(2) $L(\pi_{\theta})=-\sum_{b\in\mathcal{B}}\min\left(\pi^{\prime}(\bm{a}|\bm{s}),\text{clip}\left(\pi^{\prime}(\bm{a}|\bm{s}),1-\epsilon,1+\epsilon\right)\right)A^{\pi_{\theta}}(\bm{a},\bm{s}),$
where $\pi^{\prime}(\bm{a}|\bm{s})=\frac{\pi_{\theta}(\bm{a}|\bm{s})}{\pi_{\theta_{old}}(\bm{a}|\bm{s})}$, $\text{clip}(x,l,r)=\max(\min(x,r),l)$ and $\epsilon$ is the clip hyperparameter. $\pi_{\theta_{old}}$ is the actor network after the last update and remains unchanged when the current actor network $\pi_{\theta}$ is updated.
Critic update. To update the critic network, we use the MSE of the temporal-difference error (TD-error) as the loss function:
(3) $L^{TD}(V_{\phi})=\sum_{b\in\mathcal{B}}(\mathcal{R}_{t}+\gamma V_{\phi}(\bm{s}^{t+1})-V_{\phi}(\bm{s}^{t}))^{2}.$
We determine whether our algorithm has converged based on the policy loss $L(\pi_{\theta})$. Specifically, we regard the algorithm as converged when the policy loss remains nearly flat, defined as a decrease of less than 1% over 10 consecutive episodes.
4.2.3. Trust prediction
We use maximum likelihood estimation (MLE) to predict creators’ trust degrees. The training dataset $\mathcal{D}=\{(i,y)\}$ is collected from the interactions between the platform and creators, where $i$ denotes the creator’s ID, and $y\in\{0,1\}$ is a binary label indicating whether creator $i$ follows the platform’s suggestion. A neural network parameterized by $\psi$, denoted as $P_{\psi}$, is employed to model the creators’ trust. The network takes a creator’s ID as input and outputs the estimated trust degree, i.e., $\hat{d}_{i}=P_{\psi}(i)$. The parameters $\psi$ are learned by minimizing the MLE loss:
$L(P_{\psi})=-\sum_{(i,y)\in\mathcal{D}}\left[y\log(P_{\psi}(i))+(1-y)\log(1-P_{\psi}(i))\right].$
Figure 2. Comparison of user clicks in RS between our method and other baselines on two different datasets.
5. Experiments
In this section, we conduct extensive experiments to verify the effectiveness of our method in improving the long-term welfare of RS by addressing the following research questions:
RQ1: Can our method improve the long-term welfare of RS?
RQ2: What is the inherent insight that our method can enchance the long-term welfare?
RQ3: How robust our method is, and whether it can perform well under various settings?
The source code of LoRe and datasets are available  [^3] https://anonymous.4open.science/r/LoRe-E792.
5.1. Experimental Setups
Real-world dataset. We conduct experiments on two large-scale recommendation datasets, both collected from real-world recommendation platforms. YouTube is a dataset crawled from the world’s leading recommendation platform (Ye et al., 2025). It contains rich information about content creators across diverse categories, along with user comments on their uploaded items. The dataset comprises 4,004 creators, 1.97 million users, and 0.19 million distinct items spanning 14 genres. Amazon Video Games (Amazon)  [^4] https://amazon-reviews-2023.github.io/ is a widely used recommendation dataset. It consists of user-generated reviews, ratings, and product metadata collected from the Amazon e-commerce platform. It contains 2.8M users, 137.2K items, and 1326 creators.
Simulation setups. To better approximate real-world RS environments, we adopt the multi-stakeholder RS simulation platform proposed by Ye et al. (2025), which integrates LLM-based agents to simulate the dynamic behaviors of both users and creators. Recent studies have demonstrated that LLMs are capable of accurately mimicking user preferences and the decision-making of bounded-rational creators (Wang et al., 2024; Zhang et al., 2024; Wang et al., 2025; Ye et al., 2025). To test the robustness and generalization ability of our method under diverse environments, we also conduct additional experiments under environments where creators are simulated by non-LLM-based models (See Section 5.4). Besides, we consider the active behaviors of creators and users, as well as creator retention, following previous works (Mladenov et al., 2020; Xu et al., 2023; Ye et al., 2025): each creator $c_{i}\in\mathcal{C}$ and each user $u_{i}\in\mathcal{U}$ has probability $p_{c_{i}}$ and $p_{u_{i}}$ to be active, respectively. In each round, these probabilities determine whether a creator or user is active. Only active creators create items, and only active users receive recommendations. Moreover, a creator $c_{i}$ is assumed to leave the platform if no clicks are received in $l$ consecutive rounds of creation.
Table 2. The content diversity (CD) and average number (AN) of active creators across different methods on different datasets.
YouTube Amazon
Method CD AN CD AN
LoRe (ours) 3.259 16.250 3.462 14.710
Reweight 3.223 15.625 2.744 10.530
P-MMF 3.196 15.930 2.697 11.025
FairRec 3.182 14.245 2.646 9.880
Most History Click 3.197 14.730 2.805 11.610
Most Click 3.186 13.840 2.847 11.775
w/o signal 3.177 11.750 2.546 9.665
Baselines. We compared our method with several baselines, including: w/o Signal: revealing no information to the creator at all, except for their own creation history; and some naive information revelation strategies: Most History Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres the creator has ever created; Most Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres. In addition, we compared our method with approaches that focus on recommendation algorithm design, such as P-MMF (Xu et al., 2023) and FairRec (Guo et al., 2023). We also included a baseline that influences creators by adjusting their rewards, namely Reweight (Yao et al., 2024).
Metrics. Following previous works (Ye et al., 2025; Yao et al., 2024), we use the cumulative user click number to measure the long-term welfare. We adopt RecAgent (Wang et al., 2025) as the user simulator, where users can choose to click or skip when presented with an item. The number of clicks by all users in each round is recorded as the user welfare for that round.
Figure 3. The number of genres created by each creator.
5.2. Effectiveness of LoRe (RQ1)
To evaluate the effectiveness of LoRe in enhancing long-term user welfare, we compare its performance against several baselines on two recommendation datasets. In this experimental setting, the number of creators is fixed at 50 and the number of users at 100. The platform employs DIN (Zhou et al., 2018) as the based recommendation model, and both user and creator agents are powered by LLaMA3-8B (Touvron et al., 2023). In addition, we initialize creators’ trust degrees using a truncated Gaussian distribution, reflecting real-world scenarios where agents exhibit moderate trust levels, and extremely high or low trust is uncommon (Teacy, 2006; Teacy et al., 2012). Denoted by $\mathcal{N}(\mu,\sigma;0,1)$ a truncated Gaussian distribution with mean $\mu$, variance $\sigma$, lower bound 0, and upper bound $1$. Then, each creator’s trust degree is randomly sampled from distribution $\mathcal{N}(0.5,1.0;0,1)$.
Figure 2 shows the changes in cumulative user clicks per round as the number of rounds increases. LoRe consistently achieves higher user welfare than all baselines on both the YouTube and Amazon datasets. Compared with approaches that only adjust recommendation algorithms or creator rewards (e.g., PMMF, FairRec, Reweight), LoRe performs substantially better, indicating that influencing creator behavior through selective information revelation is more effective for improving long-term user welfare.
Existing methods also show inconsistent performance across datasets. For instance, Reweight outperforms all other baselines on YouTube but performs worse than naive information revelation strategies on Amazon. This is primarily due to the more skewed user distribution in the Amazon dataset compared to the YouTube dataset, i.e., the majority of users in the Amazon dataset favor a particular genre of content, while fewer users favor other genres, which means that existing methods may significantly degrade in performance under such conditions.
Naive information revelation strategies (Most History Click, Most Click) also display unstable performance: they outperform w/o Signal on Amazon but are similar to it on YouTube. This suggests that naive strategies lack robustness across recommendation environments. In contrast, LoRe consistently achieves the best long-term welfare, demonstrating that strategic information revelation is not only more effective than naive approaches but also more robust across different environments.
Figure 4. Training time cost and the influence of trust degree.
5.3. Inherent insight of LoRe (RQ2)
To better understand the inherent insight of LoRe, we analyze the dynamics of RS from both macro and micro perspectives. At the macro level, we examine the content diversity and the average number of active creators during the interactions among the platform, creators, and users. At the micro level, we investigate the variation in the number of content genres produced by individual creators under different methods.
Figure 5. The performance of different methods under various creator simulation models.
Macro perspective. To evaluation the content diversity, we employ entropy which is a widely used metric for measuring diversity (Spellerberg and Fedor, 2003). Let $p$ denote the probability density function over genre set $\mathcal{G}$, where $p(g_{i})\ (g_{i}\in\mathcal{G})$ represents the probability of content genre $g_{i}$. Then, the content diversity is defined as $D(p)=\sum_{g_{i}\in\mathcal{G}}-p(g_{i})\log p(g_{i})$. Diversity is an important property of RS, and increasing diversity is often associated with improved user welfare (Castells et al., 2021). Intuitively, higher diversity reflects a more balanced distribution of content genres, whereas lower diversity corresponds to a skewed content distribution in which most content is concentrated in only a few categories. Such extreme content distribution may prevent some users from accessing their preferred content, resulting in user churn and reduced user welfare. Therefore, greater diversity is more conducive to improving users’ long-term welfare. As shown in Table 2, our method achieves the highest content diversity across different datasets. This demonstrates that our approach enhances the diversity of RS compared with other baselines, which in turn indirectly improves user welfare.
In our simulations, we assumed that a creator would exit after receiving no clicks for 10 consecutive rounds. We calculated the average number of active creators under different methods across datasets. As shown in Table 2, our method retains more creators than other baselines. This improvement arises because the platform can influence creators’ behavior by suggesting them creating specific content genre, thereby guiding them to produce content that attracts more clicks.
Micro perspective. We further analyze the variation in the number of content genres produced by each creator. As shown in Figure 3, the x-axis represents the creator ID, and the y-axis indicates the number of distinct genres created by each individual. Obviously, compared to the baseline "w/o Signal", our approach leads to an increase in the number of genres per creator, which directly contributes to the overall improvement in content diversity.
5.4. Robustness Analysis (RQ3)
In this section, we conduct extensive experiments to evaluate the robustness of our method.
Training cost. We evaluate the convergence rounds and net training time of our algorithm at different scales to demonstrate its scalability in large-scale simulations. As shown in Figure 4 (a), our algorithm requires only a few rounds to converge. When the number of agents is 50, it converges in fewer than 70 rounds; when the number of agents increases to 1,000, convergence is achieved in fewer than 40 rounds. In addition, the net training time per round remains within an acceptable range. Even with 1,000 agents, the net training time per round is less than 4 seconds. These results confirm that our method scales effectively to large-scale, real-world scenarios.
Influence of creator simulation models. As discussed in Section 3.3, our method is capable of handling various creator simulation models. To verify this and evaluate its robustness, we conduct experiments across different creator simulation models. Specifically, we consider both LLM-based simulation models: CreAgent (Ye et al., 2025) (with Mistral-7B and Qwen2.5-7B as backbones), and heuristic-based models: Creator Feature Dynamics (CFD) (Lin et al., 2024; Zhan et al., 2021), where creators adjust their creation strategies using user feedback as gradients scaled by a learning rate, and SimuLine (Zhang et al., 2023), where creators’ subsequent creations are determined by probabilistic sampling based on the number of clicks in previous steps.
Figure 5 presents the performance comparison between our method and the baselines under different creator simulation models. The results clearly show that, regardless of whether the simulation model is LLM-based or heuristic-based, our approach consistently outperforms all baselines. In contrast, existing recommendation-algorithm-based methods (i.e., P-MMF and FairRec) exhibit inconsistent performance across different creator simulation models. For instance, FairRec performs best among baselines under the Mistral-7B, Qwen2.5-7B, and SimuLine simulation models but is outperformed by P-MMF under the CFD model. Moreover, Reweight, which influences creator behavior by adjusting their rewards, performs relatively poorly. This is because it is designed for a specific creator decision model and thus may fail to generalize well when the decision model changes.
Table 3. The average user clicks achieved through combining LoRe with FairRec or P-MMF on YouTube and Amazon datasets.
Method YouTube Amazon
LoRe (ours) 137.4954 95.1499
P-MMF 108.5688 51.2752
FairRec 113.2789 46.1865
LoRe + P-MMF 141.2263 101.9969
LoRe + FairRec 146.0092 98.6177
Influence of creators’ trust degree. Although platforms can provide various suggestions to creators, real-world creators differ in their degree of trust toward the platform. Highly trusting creators are more likely to follow the platform’s suggestions, whereas those with low trust are less likely to be influenced. In addition, in practice, creators’ trust levels may evolve over time according to the number of clicks they receive Prasad et al. (2023).
To evaluate the effectiveness of our method under different levels of creator trust, we conduct experiments where creators’ trust degrees are sampled from distributions with varying expectations. Specifically, we conduct experiments with $\mu\in\{0.4,0.6,0.8\}$ and $\sigma=1.0$. Moreover, we assume that creators’ trust degrees change dynamically over time. Let $d_{i}^{t}$ denote creator $i$’s trust degree in round $t$ and $r_{i}^{t}$ represent the number of clicks received by creator $i$ in round $t$. Then, creator $i$’s trust in round $t+1$ is updated as $d_{i}^{t+1}=d_{i}^{t}+\frac{r_{i}^{t+1}-r_{i}^{t}}{r_{i}^{t}}$, if they follow the platform’s suggestion. This means that a creator’s trust increases if following the platform’s suggestion yields more clicks, and decreases otherwise.
Figure 4 (b) shows the user clicks under trust degree distributions with different expectations on the YouTube dataset. We compare our method with Reweight, which performs best among the baselines on YouTube. The results show that although our method’s performance fluctuates slightly across different trust distributions, it consistently outperforms Reweight, demonstrating robustness to varying levels of creator trust. Furthermore, while performance of our method decreases slightly as the expected trust degree decreases, the decline is marginal. This trend is reasonable, as lower trust leads to fewer creators adopting the platform’s suggestions. These findings highlight the importance of platforms maintaining their credibility by providing beneficial suggestions to creators.
Incorporated with existing methods. We integrate our method with existing approaches that focus on recommendation algorithm design, such as P-MMF and FairRec. We exclude Reweight from this combination, as both LoRe and Reweight aim to enhance user welfare by influencing creators’ behavior. Table 3 presents the average user clicks achieved by combining our method with other approaches on the YouTube and Amazon datasets. It is evident that combining our method with these approaches further improves the performance. Moreover, LoRe + FairRec achieves better performance on the YouTube dataset, whereas LoRe + P-MMF performs better on the Amazon dataset. These results demonstrates that our approach is compatible with other methods and highlights its scalability.
6. Conclusions
In this paper, we reveal that influencing creators’ behaviors could be a better way to improve user welfare. Inspired by the idea of Bayesian persuasion, we model the platform as the sender and the creators as the receivers. To overcome the problem of bounded rationality assumption, we model the creators as bounded rational agents. Then we propose a novel method, LoRe, which aims to learn how to influence creators by revealing information to them. Lore introduces a reinforcement learning algorithm to address the limitations imposed by the strong assumption in Bayesian persuasion. Extensive experiments on two real-world datasets demonstrate the effectiveness of our approach and show that it can be seamlessly integrated into the existing RS ecosystem, as well as combined with other methods that focus on recommendation algorithm design.
Appendix
Appendix A Examples of Content Suggestion in Real Applications
Figure 6. Examples of creation suggestions in real-world applications across different short video platforms, each of which has hundreds of millions of users.
As the number of creators on short-video platforms continues to grow, providing creation suggestions to guide creators, thereby indirectly improving user experience, has become a key component of platform operations. Figure 6 presents examples of creation suggestions on short-video content platforms, each of which typically serves hundreds of millions of users.
Although current commercial platforms have begun to adopt creation suggestions, they are typically limited to simple suggestions, such as suggesting visual effects or the most trending topics. Effectively guiding creators in real-world industrial settings remains an open challenge. To bridge this gap, we aims to explore how to design more effective suggestion strategies in real-world platform environments to enhance overall user welfare.
Appendix B Omitted Definitions in Section 4
We respectively define the state value function $V^{\pi}(\bm{s}^{t})$, the state-action value function $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$as:
$V^{\pi}(\bm{s}^{t})=\operatorname{\mathbb{E}}_{(\bm{a}^{t},\bm{s}^{t+1},\dots)\sim\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right],$
$Q^{\pi}(\bm{s}^{t},\bm{a}^{t})=\operatorname{\mathbb{E}}_{(\bm{s}^{t+1},\bm{a}^{t+1},\dots)\sim\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right].$
Appendix C More Detailed Experimental Setups
Our experimental server is equipped with 28 Intel(R) Core(TM) i9-10940X CPUs, 128GB of RAM, and 4 NVIDIA RTX 4090 GPUs, providing a total of 96 GB of VRAM.
Appendix D Additional Experiments
D.1. Influence of Different Recommendation Models
Figure 7. Comparison of user welfare with and without our method under the RS environment with two different recommendation models, i.e, MF and BPR, respectively.
In real-world RS environments, platforms adopt different recommendation algorithms to select content from the item pool for users. To verify the effectiveness of our method under various recommendation models, we conduct experiments using different recommendation algorithms as the platform’s policy. We select two of the most widely used recommendation models: Matrix Factorization (MF) (Koren et al., 2009) and Bayesian Personalized Ranking (BPR) (Rendle et al., 2012).
As illustrated in Figure 7, across different recommendation models, our method consistently achieves better user welfare compared to the w/o signal baseline. Specifically, when using MF as the recommendation model, the user welfare with information signaling is 1.18 times that of the no-signal case (at the last round of simulation). When using BPR, this number increases to 1.26 times.
Appendix E Limitations
In our work, we use large language models (LLMs) to simulate the behaviors of users and content creators. Although previous work has verified the effectiveness of the LLM-empowered simulation agents in simulating users click behavior and creator content creation behavior, there is limited research on whether such agents can effectively simulate creators’ responses to platform-provided suggestions. Therefore, how to simulate the behavior of real-world creators after receiving platform suggestions is also a key challenge and represents a promising direction for future research.
Appendix F Social Impacts
Our method can effectively enhance the long-term user welfare in RS. This not only improves user satisfaction and engagement over time but also contributes to the development of a healthier and more sustainable ecosystem for content consumption and creation. To the best of our knowledge, our method does not pose any adverse effects on society.

### v2

@@
-Towards Long-Term User Welfare in Recommender Systems via Creator-Oriented Information Revelation
+Enhancing Long-Term Welfare in Recommender Systems: An Information Revelation Approach
 Abstract.
-Improving the long-term user welfare (e.g., sustained user engagement) has become a central objective of recommender systems (RS).
-In real-world platforms, the creation behaviors of content creators plays a crucial role in shaping long-term welfare beyond short-term recommendation accuracy, making the effective steering of creator behavior essential to foster a healthier RS ecosystem.
-Existing works typically rely on re-ranking algorithms that heuristically adjust item exposure to steer creators’ behavior.
-However, when embedded within recommendation pipelines, such a strategy often conflicts with the short-term objective of improving recommendation accuracy, leading to performance degradation and suboptimal long-term welfare.
-The well-established economics studies offer us valuable insights for an alternative approach without relying on recommendation algorithmic design: revealing information from an information-rich party (sender) to a less-informed party (receiver) can effectively change the receiver’s beliefs and steer their behavior.
+Improving long-term user welfare (e.g., sustained user engagement) has become a central objective in recommender systems (RS).
+In real-world platforms, the creation behaviors of content creators play a crucial role in shaping long-term welfare beyond short-term recommendation accuracy, making effective steering of creator behavior essential for fostering a healthy RS ecosystem.
+Existing approaches typically rely on re-ranking algorithms that heuristically adjust item exposure to influence creators’ behavior.
+However, when integrated into recommendation pipelines, such methods often conflict with the short-term objective of improving recommendation accuracy, leading to performance degradation and suboptimal welfare.
+The well-established studies in economics provide valuable insights into an alternative approach without relying on recommendation algorithm design: revealing information from an information-rich party (sender) to a less-informed party (receiver) can effectively influence the receiver’s beliefs and steer their behavior.
 Inspired by this idea, we propose an information-revealing framework, named Long-term Welfare Optimization via Information Revelation (LoRe).
-In this framework, we utilize a classical information revelation method (i.e., Bayesian persuasion) to map the stakeholders in RS, treating the platform as the sender and creators as the receivers.
-To address the challenge posed by the unrealistic assumption of traditional economic methods, we formulate the process of information revelation as a Markov Decision Process (MDP) and propose a learning algorithm trained and inferred in environments with boundedly rational creators.
-Extensive experiments on two real-world RS datasets and across different creator modeling approaches demonstrate that our method can effectively outperform existing fair re-ranking methods and information revealing strategies in improving long-term user welfare.
+In LoRe, we adopt a classical information revelation method (i.e., cheap talk) to map the stakeholders in RS, treating the platform as the sender and creators as the receivers.
+To address the challenge posed by the unrealistic assumption of traditional economic method, we formulate the process of information revelation as a Markov Decision Process (MDP) and develop a learning algorithm to optimize the information revelation strategy.
+Extensive experiments on two real-world RS datasets and across different creator modeling approaches demonstrate that our method can effectively outperform existing re-ranking methods and reward-based creator-influencing method in improving long-term user welfare.
 Recommender System, Information Asymmetry, Information Revelation, Bounded Rationality
@@
 1. Introduction
-With the rapid growth of user-generated content (UGC) on the Web (Krumm et al., 2008), modern recommender platforms (e.g., TikTok, YouTube) have evolved into complex ecosystems where users and content creators continuously interact and influence each other (Bobadilla et al., 2013; Ye et al., 2025).
-In such dynamic environments, platforms are now shifting their focus from solely optimizing short-term objectives, such as maximizing users’ immediate engagement behaviors (e.g., clicks or purchases), to fostering long-term user welfare (i.e., sustained user engagement), aiming to ensure the lasting health and sustainability of the ecosystem (Xu et al., 2023; Guo et al., 2023).
-To enhance the long-term user welfare of recommender systems (RS), previous works have shown that the creators’ creation behavior (i.e., creating new items) is a key factor (Ye et al., 2025; Xu et al., 2023; Yao et al., 2024) beyond recommendation accuracy (i.e., prediction accuracy of user behavior).
-This is because creators continuously upload new items based on previous user feedback, which shapes the item pool of RS and indirectly influences what users are exposed to in the long run (Ben-Porat and Tennenholtz, 2018; Ben-Porat et al., 2019; Boutilier et al., 2024).
-Considering the importance of creators’ behavior, most existing works utilize heuristic re-ranking algorithms to steer creators’ content creation.
-For example, they increase item exposure for small creators to enhance creator retention and stimulate content creation (Guo et al., 2023; Xu et al., 2023; Ye et al., 2024).
-However, previous works have shown that such a re-ranking strategy, when embedded within recommendation pipelines, often conflicts with the short-term objectives of improving accuracy, inevitably trading off recommendation performance and harming users’ short-term experience (Ye et al., 2024; Xu et al., 2023).
-Therefore, how to steer creators’ content creation behavior beyond relying on recommendation algorithmic design to enhance the long-term welfare remains a significant challenge in RS.
+With the rapid growth of user-generated content (UGC) (Krumm et al., 2008), modern recommender platforms (e.g., TikTok, YouTube) have evolved into complex ecosystems where users and content creators continuously interact and influence each other (Bobadilla et al., 2013; Ye et al., 2025).
+In such dynamic environments, platforms are now shifting their focus from solely optimizing short-term objectives, such as maximizing users’ immediate engagement behaviors (e.g., clicks), to fostering long-term user welfare (i.e., sustained user engagement), aiming to ensure the lasting health and sustainability of the platform ecosystem (Xu et al., 2023; Guo et al., 2023).
+To enhance the long-term user welfare of recommender systems (RS), previous studies have shown that the creators’ creation behavior (i.e., creating new items) is a key factor (Ye et al., 2025; Xu et al., 2023; Yao et al., 2024) beyond recommendation accuracy (i.e., prediction accuracy of user behavior).
+This is because creators continuously upload new items based on previous user feedback, thereby shaping the item pool of RS and indirectly influences what users are exposed to in the long run (Ben-Porat and Tennenholtz, 2018; Ben-Porat et al., 2019; Boutilier et al., 2024).
+Given the importance of creator behavior, most existing works utilize heuristic re-ranking algorithms to steer creators’ content creation.
+For example, they increase item exposure for smaller creators to enhance creator retention and stimulate content creation (Guo et al., 2023; Xu et al., 2023; Ye et al., 2024).
+However, prior work have shown that such re-ranking strategies, when embedded within recommendation pipelines, often conflict with the short-term objectives of improving accuracy, inevitably trading off recommendation performance and harming users’ short-term experience (Ye et al., 2024; Xu et al., 2023).
+Consequently, how to steer creators’ content creation behavior without relying on recommendation algorithmic design to enhance the long-term welfare remains a significant challenge in RS.
 The investigation of behavioral steering among strategic agents has a long history in economics (Milgrom and Roberts, 1987; Dierkens, 1991; Mishra et al., 1998; Crawford et al., 2018), offering valuable insights for enhancing the long-term welfare of RS.
-In economics studies, a common approach is to leverage information revelation under conditions of information asymmetry (i.e., agents possess unequal amounts of information) to influence agents’ behavior.
-Such a strategy aims to design an information revelation strategy, which specifies what information the information-rich party (sender) should disclose to the less-informed party (receiver).
-By revealing certain information unavailable to the receiver, the receiver can update their beliefs and adjust their strategy accordingly, leading to observable behavioral changes.
+In economics studies, a common approach is to leverage information revelation under the situation of information asymmetry (i.e., agents possess unequal amounts of information) to influence agents’ behavior.
+Such a strategy focuses on designing an information revelation strategy that determines what information an information-rich party (sender) should disclose to a less-informed party (receiver).
+By revealing certain information unavailable to the receiver, the receiver can update the beliefs and adjust the action accordingly, resulting in observable behavioral changes.
+The revealed information in such situation can be generally interpreted as a suggestion that guides the receiver’s action.
 Inspired by this idea, we can steer creators’ behavior to enhance long-term welfare by revealing information, viewing the platform as the sender and creators as the receivers.
@@
 Similar but simpler idea have been adopted in industrial practice, offering real-world scenarios and potential applications for more advanced information revelation approaches.
-For example, platforms like TikTok provide creators with suggestions to produce trending content (see Appendix A).
+For instance, platforms like TikTok and YouTube provide creators with suggestions to produce trending content.
 However, directly applying information revelation to RS is challenging.
-Traditional methods in economics studies (Yao et al., 2024; Prasad et al., 2023; Boutilier et al., 2024) typically assume that receivers are completely rational, i.e., they can always select the optimal strategy to maximize their utility given the available information.
-In contrast, real-world creators are bounded rational (Simon, 1990; Kahneman and Tversky, 2013), because their decision-making is constrained by the limited cognitive capacity, attention, and information-processing ability of humans.
-To address this challenge, we propose our information-revealing framework tailored for RS, named Long-term Welfare Optimization via Information Revelation (LoRe).
-Specifically, we map the key components of classical information revelation methods in economics (i.e., Bayesian persuasion (Gan et al., 2022a; Kamenica and Gentzkow, 2011)) onto the corresponding stakeholders in RS.
-Based on this mapping, we formulate the platform’s information revelation process as a Markov Decision Process (MDP) (Puterman, 1990) and propose a learning algorithm to optimize the information revelation strategy.
-Instead of relying on the assumption of completely rational agents, LoRe is trained and inferred within a realistic reinforcement learning (RL) environment, where bounded rational users and creators interact with each other.
+Traditional economic approaches (Yao et al., 2024; Prasad et al., 2023; Boutilier et al., 2024) typically assume that receivers are completely rational, i.e., they always select the optimal strategy to maximize their utility.
+Under this assumption, receivers’ behavior can be computed explicitly, enabling the sender to optimally send information to maximize its own utility.
+In contrast, real-world creators are boundedly rational (Simon, 1990; Kahneman and Tversky, 2013), because their decision-making is constrained by the limited cognitive capacity, attention, and information-processing ability of humans.
+Consequently, it is difficult to calculate creators’ behavior in real-world precisely, preventing the platform from directly computing the optimal information to reveal.
+To address this challenge, we propose an information revelation framework tailored for RS, named Long-term Welfare Optimization via Information Revelation (LoRe).
+Specifically, we map the key components of classical information revelation method in economics studies (i.e., cheap talk model (Farrell and Rabin, 1996)) onto the corresponding stakeholders in RS.
+Based on this mapping, we formulate the platform’s information revelation process as a Markov Decision Process (MDP) (Puterman, 1990) and develop a learning algorithm to optimize the platform’s information revelation strategy.
 Unlike previous works that rely on heuristic exposure allocation in the re-ranking stage to steer creators’ behavior (Guo et al., 2023), LoRe serves as an independent module for the platform and does not depend on any specific recommendation algorithm.
-Experiments on two real-world recommendation datasets demonstrate that LoRe outperforms both creator-fair re-ranking methods and simple information-revealing strategies in improving long-term user welfare (i.e., sustained user clicks).
+Experiments on two real-world recommendation datasets demonstrate that LoRe outperforms both creator-fair re-ranking methods and reward-based creator-influencing method in improving long-term user welfare (i.e., sustained user clicks).
 Moreover, since LoRe operates independently, it can be seamlessly integrated with existing ranking or re-ranking algorithms to further enhance long-term welfare.
+In our experiments, to more faithfully simulate the real-world RS environments, we adopt the multi-stakeholder simulation framework proposed by Ye et al.
+(2025), which leverages LLM-based agents to simulate the dynamic behaviors of users and creators.
+Specifically, these agents simulate user click-through behavior and creator content generation.
+This approach is grounded in recent research which demonstrate that agents empowered by LLMs can effectively emulate user preferences and the creation behavior of bounded rational creators (Wang et al., 2024; Zhang et al., 2024; Wang et al., 2025; Ye et al., 2025).
 We summarize our contributions as follows:
-(1) We highlight that steering creators’ behavior through information revealing provides an effective approach to enhancing long-term user welfare beyond relying on recommendation algorithmic design.
-(2) Inspired by economic studies, we propose a novel information revelation method, LoRe, which models the revelation process as a Bayesian persuasion problem and learns to optimize the revelation strategy in environments where boundedly rational agents interact.
-(3) Extensive experiments on real-world datasets demonstrate that our method can effectively improve the long-term user welfare of RS by steering creators’ content creation.
+(1) We highlight that influencing creators’ creation behavior through information revelation provides an effective approach to enhancing long-term user welfare beyond relying on recommendation algorithm design.
+(2) Inspired by economic studies, we propose a novel information revelation method, LoRe, which models the revelation process as a MDP and learns to optimize the information revelation strategy in environments where boundedly rational agents interact.
+(3) Extensive experiments on real-world datasets demonstrate that our method can effectively improve the long-term user welfare of RS by steering creators’ creation behavior.
 2. Related Work
@@
 (2025) introduces an LLM-based creator simulator, CreAgent, aiming to achieve a more realistic evaluation of multi-stakeholder RS.
-However, existing works mainly focus on developing and evaluating RS algorithms, and do not consider another crucial side, i.e., steering creators via creation suggestions to indirectly optimize the long-term welfare of RS.
-Information asymmetry in RS.
+However, existing works mainly focus on developing and evaluating RS algorithms, and do not consider another crucial side, i.e., steering creators via information revelation to indirectly optimize the long-term welfare of RS.
+Information asymmetry in RS and information revelation.
 A relevant topic to our work is the literature considering the information asymmetry in RS (Prasad et al., 2023; Yao et al., 2023, 2024; Boutilier et al., 2024; Ye et al., 2025).
@@
 (2024); Ye et al.
-(2025) highlight the significant information asymmetry in RS, but they do not proposed a method to improve the welfare of RS under such asymmetry.
+(2025) highlight the significant information asymmetry in RS, but they do not proposed a method to address such a problem.
 Prasad et al.
-(2023) propose a prompting policy that promises to provide creators a certain level of user clicks to influence the their content creation.
+(2023) propose a prompting policy that promises to provide creators a certain level of user clicks to influence the their creation behavior.
 Yao et al.
 (2023, 2024) suggest influencing content creation by adjusting reward weights to encourage different types of content.
-However, these studies model creators as completely rational agents, ignoring the bounded rationality observed in real-world creators, which limits the applicability and performance of their methods.
-Information revelation under information asymmetry.
-Our work is also related to the literature on information revelation under information asymmetry in economics, where a prominent approach is Bayesian persuasion (Kamenica and Gentzkow, 2011).
-In Bayesian persuasion, an informed party (sender) designs an information revelation strategy to reveal the optimal information to a less informed party (receiver) (Kamenica and Gentzkow, 2011).
-The most relevant Bayesian persuasion models to our work consider one sender (the platform) and multiple receivers (the creators) (Wang, 2013; Arieli and Babichenko, 2019; Castiglioni et al., 2021; Bernasconi et al., 2023; Zhao et al., 2025), where the sender simultaneously reveals information to multiple receivers.
-These studies investigate how to design information disclosure strategies to maximize the sender’s utility in a static setting.
-There is also a stream of research on dynamic Bayesian persuasion (Gan et al., 2022b, a; Wu et al., 2022), which examines repeated interactions between a sender and a receiver.
-However, because of their strong assumption of completely rational receivers, these models cannot be directly applied to real-world RS scenarios, where creators typically exhibit bounded rationality.
+However, these methods inevitably harm creators’ realized rewards and may introduce unfairness, as they directly alter the rewards that creators would otherwise receive.
+The literature on information revelation under information asymmetry in economics is also related to our work, where a prominent approach is cheap talk (Farrell and Rabin, 1996; Teacy, 2006; Chakraborty and Harbaugh, 2010; Ivanov, 2015).
+In cheap talk, an informed party (sender) designs an information revelation strategy to reveal the information to a less informed party (receiver).
+However, because of their strong assumption of completely rational receivers, these models cannot be directly applied to real-world RS scenarios where creators typically exhibit bounded rationality.
+LLM-based simulator in RS.
+In recent years, given the superior semantic understanding and reasoning capabilities of LLMs, LLM-based simulators for recommender systems have attracted increasing attention and experienced rapid development.
+They are now widely regarded as a reliable technique for modeling user behavior in recommender systems (Ye et al., 2025; Ma et al., 2025; Zhu et al., 2025; Zhang et al., 2025).
+Initially, research (Ma et al., 2025; Zhu et al., 2025; Zhang et al., 2025) focused on simulating user behavior in RS, such as clicks, to obtain real-world user feedback for evaluating recommendation algorithms.
+Most recently, the field has progressed toward multi-stakeholder and ecosystem simulation, recognizing that RS operate within complex environments involving users, creators and platforms (Wang et al., 2024; Ye et al., 2025; Coppolillo et al., 2025).
+These studies demonstrated that LLM-based RS simulators can replicate the behaviors of both creators and users in real-world, providing a reliable framework to evaluate the performance of algorithms developed for such systems.
 3. Preliminaries
-In this section, we first introduce the settings of real-world RS. Then we introduce the basic idea of Bayesian Persuasion. Finally, we describe how to model creators as bounded rational agents.
 3.1. Recommender Systems
@@
 Specifically, we use $\bm{Y}$ to represent the click matrix, where $\bm{Y}_{i,j,t}=1$ if item $i\in\mathcal{I}$ is clicked by user $u_{j}$ in round $t$ and 0 otherwise.
-In addition, we use $\mathcal{H}_{c_{j}}=\{(i,k_{i})|i\in\mathcal{I}\}$ to denote historical information of items created by creator $c_{j}$, where $(i,k_{i})\in\mathcal{H}_{c_{j}}$ means that the item $i\in\mathcal{I}$ produced by creator $c_{j}$ has been clicked $k_{i}$ times.
+In addition, we use $\mathcal{H}_{c_{j}}=\{(i,k_{i})\mid i\in\mathcal{I}\}$ to denote historical information of items created by creator $c_{j}$, where $(i,k_{i})\in\mathcal{H}_{c_{j}}$ means that the item $i\in\mathcal{I}$ produced by creator $c_{j}$ has been clicked $k_{i}$ times.
 For easy notation, let $\mathcal{H}=\{\mathcal{H}_{c_{1}},\mathcal{H}_{c_{2}},\dots,\mathcal{H}_{c_{|\mathcal{C}|}}\}$ denote the set of all creators’ historical information.
@@
 Obviously, there is a significant information asymmetry between platform and creators.
-Therefore, how to improve the long-term welfare under information asymmetry is an unexplored and urgent problem to be addressed.
-Figure 1. The overall framework of our method LoRe, which consists two phases: Online Inference and Offline Training.
-3.2. Bayesian Persuasion
-Information asymmetry formulation.
-The standard Bayesian persuasion considers an idealized model with a sender and a receiver.
+Figure 1.
+The overall framework of our method LoRe, which consists two phases: Inference and Training.
+During the inference phase, in each round, the platform sends suggestions to creators, who then produce content, after which users decide whether to click on the recommended items.
+The interaction data generated in each round are collected by the platform.
+In the training phase, the platform uses the data collected during inference to train the actor and critic networks.
+3.2. Cheap Talk in Recommender Systems
+Cheap talk.
+The standard cheap talk model considers an idealized setting with a sender and a receiver.
 Both the receiver’s utility function $\mu_{r}(a_{r},w)$ and the sender’s utility function $\mu_{s}(a_{r},w)$ depend on receiver’s action $a_{r}\in\mathcal{A}_{r}$ and state $w\in\mathcal{W}$.
-The state $w$ follows a cumulative distribution function $F(w)$ with probability density function $f(w)$.
-It is assumed that both sender and receiver know $F(w)$, but the sender exactly knows the real state $w$ after $w$ is sampled.
-Thus, the sender have more information than the receiver, which causes the information asymmetry.
-Bayesian persuasion under information asymmetry.
-The receiver will choose an optimal action $a_{r}^{*}$ to maximize her expected utility, and the sender will also obtain a corresponding utility.
-However, the sender can use a so-called signaling scheme to reveal the receiver some information, and then influence receiver’s action.
-Specifically, we use $\sigma=(\mathcal{A}_{s},\pi)$ to denote a signaling scheme, where $\mathcal{A}_{s}$ is the set of possible signals (messages) and $\pi:\mathcal{W}\times\mathcal{A}_{s}\mapsto[0,1]$ is the signaling strategy which maps the state to a distribution of signals.
-For instance, $\pi(a_{s}|w)$ denotes the probability of sending signal $a_{s}\in\mathcal{A}_{s}$ when the real state is $w$.
-According to the revelation principle (Myerson, 1979), each signal $a_{s}$ can be viewed as an action suggestion for receiver, that is $|\mathcal{A}_{r}|=|\mathcal{A}_{s}|$, where $a_{s}\in\mathcal{A}_{s}$ means suggesting receiver taking corresponding action $a_{r}\in\mathcal{A}_{r}$ and receiver will follows sender’s suggestion [^1] Some constraints can be used to ensure that the receiver will follow the sender’s suggestion, and these constraints do not affect the optimal solution.
-For details, please refer to the reference (Kamenica and Gentzkow, 2011)..
-By designing the optimal strategy $\pi^{*}$, the sender can improve both the utility of sender and receiver under information asymmetry (Kamenica and Gentzkow, 2011).
-Table 1. Correspondence between elements in the Bayesian Persuasion and RS. See Section 4 for undefined notations.
-Bayesian Persuasion Recommender System
+In this model, only the sender exactly knows the true state $w$, whereas the receiver does not.
+Thus, the sender possesses more information than the receiver, giving rise to information asymmetry between sender and receiver.
+When the receiver has no additional information, she chooses an optimal action $a_{r}^{*}$ to maximize her expected utility based on what she knows, and the sender correspondingly obtains utility $\mu_{s}(a_{r}^{*},w,b)$.
+However, the sender can influence the receiver’s action by revealing information (message).
+Specifically, we use $M$ to denote the set of possible messages the sender can send to receiver.
+The specific form of a message is immaterial, as long as it is commonly understood and interpretable by both parties.
+Intuitively, a specific message $m\in M$ can be viewed as a suggestion of the form: "Based on the available information, I recommend that you take action $\hat{a}_{r}$".
+Formally, we use $\pi:\mathcal{W}\times M\mapsto[0,1]$ to denote the sender’s information revelation strategy, where $\pi(m\mid w)$ represents the probability of sending message $m\in M$ when the real state is $w$.
+Upon receiving a message, the receiver updates her beliefs and chooses an optimal action accordingly.
+Previous researches (Farrell and Rabin, 1996; Teacy, 2006; Chakraborty and Harbaugh, 2010; Ivanov, 2015) shows that, by appropriately designing the information revelation strategy, the sender can improve the utilities of both the sender and the receiver under information asymmetry.
+Table 1. Correspondence between elements in Cheap Talk and RS. See section 4 for undefined notations.
+Cheap Talk Recommender System
 sender platform
 receivers creators
-state $w$ content distribution $\bm{e}$ and trust $\bm{d}$
-signal set $\mathcal{A}_{s}$ suggestion set $\mathcal{A}_{p}$
-signaling strategy $\pi$ information revealing strategy $\pi$
-3.3. Bounded Rationality Modeling
-Motivated by the advantage of Bayesian persuasion under information asymmetry, we adopt the idea of this framework to guide how the platform reveals information in RS.
-Simply put, we allow the platform to send suggestions to creators, i.e., suggesting creators creating specific genres of content, which is called the platform’s information revelation strategy.
-In the standard Bayesian persuasion, receivers are assumed to be completely rational, which means they can precisely choose an action $a_{r}^{*}$ to maximize their utility $\mu_{r}(a_{r},w)$.
-This also allows senders to design signaling strategy that ensure that receivers completely follow the sender’s action suggestion.
-However, creators in reality are typically bounded rational, as they cannot compute a specific utility function to maximize their utility.
-Thus, platform in RS cannot make sure that creators will follow platform’s suggestion.
-To model creators’ bounded rationality, inspired by Prasad et al.
-(2023), we design a trust-based model, where each creator $c_{i}$ has a trust degree $d_{{}_{i}}\in[0,1]$.
-When creator $c_{i}$ receives platform’s suggestion, the probability that creator $c_{i}$ follows the suggestion is $d_{{}_{i}}$ .
-If creator $c_{i}$ dose not follow platform’s suggestion, we assume that he can use any strategy that does not perform utility calculation to choose an action, such as randomly selecting from the genres of previously created content, or selecting the most popular genre of previously created content, etc.
-That is, we do not restrict the creators’ decision model, and any decision model can be embedded in our approach that is introduced in next section.
-In the experiment section, to verify the universality of our method for different creator decision models, we use different model to simulate creators’ decision-making, such as LLM-based simulation model and heuristic-based simulation models.
-We correspond the elements of RS with those in the Bayesian persuasion, as shown in Table 1.
-$\bullet$ The platform and creators are regarded as the sender and receivers, respectively.
-This is because the platform, who possesses more information, can guide the creator’s creation by revealing information.
-$\bullet$ The content distribution in RS and creators’ trust degree are viewed as the state in Bayesian persuasion.
-$\bullet$ The suggestion that platform sends to the creators is considered as the signal in Bayesian persuasion.
-$\bullet$ The platform’s strategy of sending suggestions (information revelation strategy $\pi$) is corresponds to the signaling strategy $\pi$ in Bayesian persuasion.
+state $w$ distribution of items $s$
+message set $M$ suggestion set $\mathcal{A}_{p}$
+information revelation strategy $\pi$ suggestion strategy $\pi_{\theta}$
+Cheap talk in Recommender Systems.
+Motivated by the advantage of cheap talk under information asymmetry, we adopt this framework to guide how the platform reveals information in RS, where information asymmetry similarly exists between the platform and content creators.
+Simply put, we allow the platform to send suggestions to creators, such as "I suggest that you create an item of genre $g_{i}\in\mathcal{G}$, as this may attract more clicks or likes".
+We map the key components of RS to their counterparts in the cheap talk model, as summarized in Table 1.
+A detailed description of these correspondences is provided below.
+We model the platform and the creators as the sender and the receivers, respectively.
+This modeling choice is motivated by the fact that the platform possesses richer information and can influence creators’ creation behavior through information revelation.
+Moreover, creators’ creation behavior shapes the item pool of RS, thereby indirectly determining the content to which users are exposed over time and, consequently, affecting long-term user welfare.
+Therefore, by steering creators’ creation behavior, platforms can indirectly improve long-term user welfare.
+In the cheap talk model, the receiver’s utility depends on the actual state, which is unobserved by the receiver.
+Analogously, in RS, we represent the state as the distribution of item genres, since this distribution has a substantial impact on clicks, likes, and other performance metrics of creators’ items, and is typically unknown to individual creators.
+For example, if an increasing number of creators produce items of the same genre, competition within that genre intensifies, reducing the traffic received by each creator.
+We denote the set of possible states by $\mathcal{S}$.
+As the platform is allowed to send suggestions to creators, the messages in the cheap talk model correspond to the platform’s suggestions in RS.
+We use $\mathcal{A}_{p}=\{a_{i}\}_{i=0}^{|\mathcal{G}|}$ to denote the set of platform’s possible suggestions, where $a_{0}$ represents sending no suggestion, and $a_{i}\ (i\neq 0)$ represents the suggestion: "I recommend that you create item with genre $g_{j}$".
+Corresponding to the sender’s information revelation strategy in the cheap talk model, in RS, we use $\pi_{\theta}:\mathcal{S}\times\mathcal{A}_{p}\mapsto[0,1]$ to represent the platform’s suggestion strategy.
+Specifically, $\pi_{\theta}(a_{i}\mid s)$ is the probability that the platform sends suggestion $a_{i}$ when the state is $s\in\mathcal{S}$.
+In the traditional cheap talk model, the sender can compute an optimal information revelation strategy to maximize his own utility because the receiver is assumed to be completely rational, and thus the receiver’s behavior in response to any message can be precisely inferred.
+In contrast, in RS, creators are not completely rational and their creation behavior after receiving platform’s suggestions is generally unknown to the platform.
+As a result, the platform cannot directly compute the optimal suggestion strategy.
+Consequently, the theoretical results of the classic cheap talk model cannot be directly applied to RS, making the optimization of the platform’s suggestion strategy a challenging problem.
 4. Our Approach: LoRe
-In this section, to take into account creators’ bounded rationality and optimize platform’s information revelation strategy, we propose a novel method, LoRe.
+In this section, to address the challenge by applying the framework of cheap talk into RS and optimizing the platform’s suggestion strategy, we propose a novel method, LoRe.
 The overall algorithm workflow of LoRe is shown in Figure 1 and Algorithm 1.
-To optimize the long-term welfare of RS, we first formulate the dynamic interaction process under information asymmetry between the platform and creators as a MDP, which is described in Section 4.1.
-Then, we design a RL algorithm to optimize the platform’s information revelation strategy under such a dynamic environment, as presented in Section 4.2.
-4.1. MDP for Platform and Creators Interactions
-In our method, we use $\mathcal{A}_{p}=\{a_{i}\}_{i=0}^{|\mathcal{G}|}$ to denote the set of platform’s suggestions, where $a_{0}$ means no suggestion and $a_{i}\ (i\neq 0)$ means that platform suggests the creator to produce item with genre $g_{i}$.
-At the beginning of each round $t$, the platform will send suggestions selected from $\mathcal{A}_{p}$ to creators.
-Note that different creators may receive different suggestions.
-After receiving the suggestion, each creator will decide whether to follow the suggestion and choose to produce an item.
-Then, for each arriving user, the platform will generate $k$ items to recommend and each user can choose to click or skip these items.
-Then, the MDP for the interaction between platform and creators can be defined as a tuple $M=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where:
+We first formulate the dynamic interaction process between the platform and creators under information asymmetry as a Markov decision process (MDP), which is described in Section 4.1.
+We then design a reinforcement learning algorithm to optimize the platform’s suggestion strategy in this dynamic environment, as presented in Section 4.2.
+4.1. Interactions Between Platform and Creators
+We now formulate the dynamic interaction between the platform and creators in each round as follows.
+At the beginning of round $t$, the platform observes the state $s^{t}\in\mathcal{S}$, which represents the distribution of item genres.
+The platform then sends each creator a suggestion from suggestions set $\mathcal{A}_{p}$.
+Upon receiving a suggestion, a creator can choose to produce an item with any possible genre $g_{i}\in\mathcal{G}$.
+We do not impose any assumption of creator’s creation behavior, which indicates that any creator’s behavior model can be embedded in our method.
+After content creation and recommendation, creator $c_{i}$ obtains the number of clicks of his item as utility, denoted by $\mu_{c_{i}}^{t}$.
+Then, the cumulative utility of creator $c_{i}$ over $T$ rounds is given by:
+$\mu_{c_{i}}=\sum_{t=1}^{T}\mu_{c_{i}}^{t}.$
+The platform’s objective is to maximize the long-term user welfare, measured by the total number of user clicks, through sending appropriate suggestions.
+With a slight abuse of notation, we use $\bm{a}^{t}=(a_{1}^{t},a_{2}^{t},\dots,a_{|\mathcal{C}|}^{t})$ to denote the platform’s suggestion strategy in round $t$, where $a_{i}^{t}\in\mathcal{A}_{p}$ is the suggestion sent to creator $c_{i}$.
+Then, the platform’s problem can be formulated as follows:
+$\max_{(\bm{a}^{1},\bm{a}^{2},\dots,\bm{a}^{T})}\sum_{c\in\mathcal{C}}\mu_{c}.$
+This interaction can be abstracted as follows.
+Given a state, the platform sends suggestions to creators, after which the system transitions to a new state as a result of creators’ creation behavior, and the platform receives a reward equal to the total number of user clicks.
+Moreover, because the state is defined by the distribution of item genres, the state transition can be reasonably approximated as depending only on the current state and the platform’s actions.
+Consequently, the dynamic interaction between the platform and creators can be naturally modeled as a Markov decision process.
+We formulate this MDP between the interaction between platform and creators as a tuple $M=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where:
 $\bullet$ State $\mathcal{S}$: $\mathcal{S}$ is the set of states. Each state is denoted by
-$\bm{s}^{t}=(e_{1}^{t},e_{2}^{t},\dots,e_{|\mathcal{C}|}^{t};\hat{d}_{1}^{t},\hat{d}_{2}^{t},\dots,\hat{d}_{|\mathcal{C}|}^{t})\in\mathcal{S},$
-which contains the distribution of item genres and the predicted creators’ trust in round t1t-1, where $e_{i}^{t}$ denotes the genre of item created by creator $c_{i}$ and $\hat{d}_{i}^{t}$ represents the predicted trust of creator $i$ in round t1t-1.
-We use $\bm{e}=(e_{1},e_{2},\dots,e_{|\mathcal{C}|})$ to denote content distribution and use $\bm{\hat{d}}=(\hat{d}_{1},\hat{d}_{2},\dots,\hat{d}_{|\mathcal{C}|})$ to represent the predicted trust vector.
-In simple terms, we have $\bm{s}^{t}=(\bm{e}^{t};\bm{d}^{t})$.
+$\bm{s}^{t}=(e_{1}^{t},e_{2}^{t},\dots,e_{|\mathcal{C}|}^{t})\in\mathcal{S},$
+which is the distribution of item genres in round t1t-1, where $e_{i}^{t}$ denotes the genre of item created by creator $c_{i}$ in round t1t-1.
 $\bullet$ Action $\mathcal{A}$: $\mathcal{A}$ is the set of the platform’s actions.
@@
 $\bullet$ Discount rate $\gamma$: $\gamma\in[0,1]$ is the discount factor that quantifies the present value of future rewards.
-Note that we incorporate predicted creators’ trust degree into the state, which can help capture creators’ bounded rational behavior.
-The platform’s strategy of sending suggestion signals is defined as $\pi:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$, where $\pi(\bm{a}^{t}|\bm{s}^{t})$ is the probability of taking action $\bm{a}^{t}$ when the state at round $t$ is $\bm{s}^{t}$.
-Then, the expected discounted user clicks under strategy $\pi$ can be formulated as:
-(1) $J(\pi)=\operatorname{\mathbb{E}}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1})\right],$
-where $\tau$ is a trajectory, i.e., $\tau=(\bm{s}^{0},\bm{a}^{0},\bm{s}^{1},\bm{a}^{1},\dots)$ and $\tau\sim\pi$ is a shorthand indicating that the distribution over trajectories depends on $\pi:\bm{a}^{t}\sim\pi(\cdot|\bm{s}^{t})$, $\bm{s}^{t+1}\sim\mathcal{T}(\cdot|\bm{s}^{t},\bm{a}^{t})$.
-Furthermore, we respectively use $V^{\pi}(\bm{s}^{t})$, $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ and $A^{\pi}(\bm{s}^{t},\bm{a}^{t})$ to denote the state value function, state-action value function and the advantage function, where $V^{\pi}(\bm{s}^{t})$ is the value of state $\bm{s}^{t}$, $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ is the value of taking action $\bm{a}^{t}$ under state $\bm{s}^{t}$ and $A^{\pi}(\bm{s}^{t},\bm{a}^{t})=Q^{\pi}(\bm{s}^{t},\bm{a}^{t})-V^{\pi}(\bm{a}^{t})$ [^2] In practice, the advantage function $A^{\pi}(\bm{s}^{t},\bm{a}^{t})$ is calculated by the generalized advantage estimation (Schulman et al., 2015)..
-$V^{\pi}(s^{t})$ and $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ are common functions in MDP, and their definitions can be found in appendix B.
-4.2. Reinforcement Learning Algorithm for Platform’s Strategy
-In this section, we aim to optimize the long-term welfare in Eq(1) under the dynamic user-creator-platform interaction environment by learning an information-revealing strategy $\pi$.
+The platform’s suggestion strategy is redefined as:
+$\pi_{\theta}:\mathcal{S}\times\mathcal{A}\mapsto[0,1],$
+where $\pi_{\theta}(\bm{a}^{t}|\bm{s}^{t})$ is the probability of taking action $\bm{a}^{t}$ when the state at round $t$ is $\bm{s}^{t}$.
+Then, the expected discounted user clicks under strategy $\pi_{\theta}$ can be formulated as:
+(1) $J(\pi_{\theta})=\operatorname{\mathbb{E}}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1})\right],$
+where $\tau$ is a trajectory, i.e., $\tau=(\bm{s}^{0},\bm{a}^{0},\bm{s}^{1},\bm{a}^{1},\dots)$ and $\tau\sim\pi_{\theta}$ is a shorthand indicating that the distribution over trajectories depends on $\pi_{\theta}:\bm{a}^{t}\sim\pi_{\theta}(\cdot|\bm{s}^{t})$, $\bm{s}^{t+1}\sim\mathcal{T}(\cdot|\bm{s}^{t},\bm{a}^{t})$.
+Furthermore, we respectively use $V^{\pi_{\theta}}(\bm{s}^{t})$, $Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ and $A^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ to denote the state value function, state-action value function and the advantage function, where $V^{\pi_{\theta}}(\bm{s}^{t})$ is the value of state $\bm{s}^{t}$, $Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ is the value of taking action $\bm{a}^{t}$ under state $\bm{s}^{t}$ and $A^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})=Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})-V^{\pi_{\theta}}(\bm{a}^{t})$ [^1] In practice, the advantage function $A^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ is calculated by the generalized advantage estimation (Schulman et al., 2015)..
+$V^{\pi_{\theta}}(s^{t})$ and $Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ are common functions in MDP, and their definitions are as follows:
+$V^{\pi_{\theta}}(\bm{s}^{t})=\operatorname{\mathbb{E}}_{(\bm{a}^{t},\bm{s}^{t+1},\dots)\sim\pi_{\theta}}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right],$
+$Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})=\operatorname{\mathbb{E}}_{(\bm{s}^{t+1},\bm{a}^{t+1},\dots)\sim\pi_{\theta}}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right].$
+4.2. Reinforcement Learning Algorithm for Platform’s Suggestion Strategy
+In this section, we aim to optimize the long-term welfare in Eq(1) under the dynamic user-creator-platform interaction environment by learning an suggestion strategy $\pi_{\theta}$.
 To achieve this, we design an actor-critic based RL algorithm.
-The overall workflow of our algorithm is shown in Algorithm 1, which contains two phases: the online inference (Line 4-13) and the offline training (Line 14-19).
+The overall workflow of our algorithm is shown in Algorithm 1, which contains two phases: the inference phase (Line 4-13) and the training phase (Line 14-19).
 0: Discount rate $\gamma$, clip hyperparameter $\epsilon$, learning rate lrlr, total rounds $T$, training epochs $M$, cumulative data rounds NN, reply buffer $\mathcal{B}=\{(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})\}$.
@@
 3:  Clear reply buffer: $\mathcal{B}=\{\}$.
-4:  // $\texttt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} Phase 1: Online Inference}}$.
+4:  // $\texttt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} Inference Phase}}$.
 5:  for $r=1$ to NN do
-6:   Get item genres’ distribution $\bm{e}^{t}$ and predict trust $\hat{\bm{d}}^{t}$.
-7:   Get state $\bm{s}^{t}=(\bm{e}^{t},\hat{\bm{d}}^{t})$.
-8:   Sample action $\bm{a}^{t}\sim\pi_{\theta}(\bm{s}^{t})$. $\texttt{Platform sends suggestions}$.
-9:   $\texttt{Creators produce items}$.
-10:   The next state transitions to $\bm{s}^{t+1}$.
-11:   $\texttt{Users interact with items (click or skip)}$.
-12:   Get the total users’ clicks $\mathcal{R}$.
-13:   Add $(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ to reply buffer $\mathcal{B}$.
-14:  end for
-15:  // $\texttt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} Phase 2: Offline Training}}$.
-16:  for $i=1$ to $M$ do
-17:   Calculate actor loss $L$ and critic loss $L^{TD}$ based on $\mathcal{B}$ according to Eq. (2) and Eq. (3).
-18:   Update actor network using the gradient $\nabla_{\pi_{\theta}}L$.
-19:   Update critic network using the gradient $\nabla_{V_{\phi}}L^{TD}$.
+6:   Get state $\bm{s}^{t}$.
+7:   Sample action $\bm{a}^{t}\sim\pi_{\theta}(\bm{s}^{t})$. $\texttt{Platform sends suggestions}$.
+8:   $\texttt{Creators produce items}$.
+9:   The next state transitions to $\bm{s}^{t+1}$.
+10:   $\texttt{Users interact with items (click or skip)}$.
+11:   Get the total users’ clicks $\mathcal{R}$.
+12:   Add $(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ to reply buffer $\mathcal{B}$.
+13:  end for
+14:  // $\texttt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} Training Phase}}$.
+15:  for $i=1$ to $M$ do
+16:   Calculate actor loss $L$ and critic loss $L^{TD}$ based on $\mathcal{B}$ according to Eq. (2) and Eq. (3).
+17:   Update actor network using the gradient $\nabla_{\pi_{\theta}}L$.
+18:   Update critic network using the gradient $\nabla_{V_{\phi}}L^{TD}$.
+19:  end for
 20:end for
-21:end for
 Algorithm 1 The overall workflow of our method LoRe.
-4.2.1. Online Inference
-At the beginning, the platform gets the current item genres’ distribution and predicts creators’ trust degree, which together constitute the state $\bm{s}^{t}$.
-We show how to predict creators’ trust degree in the following subsection.
-Then, the actor network $\pi_{\theta}$, which is parameterized by $\theta$, takes $\bm{s}^{t}$ as input and outputs the action distribution $\bm{P}^{t}=\pi_{\theta}(\bm{s}^{t})$, where the input $\bm{s}^{t}$ and output $\bm{P}^{t}$ are illustrated as follows.
+4.2.1. Inference Phase
+At the beginning, the platform gets the current item genres’ distribution, which constitute the state $\bm{s}^{t}$.
+Then, the actor network $\pi_{\theta}$, which is parameterized by $\theta$, takes $\bm{s}^{t}$ as input and outputs the action distribution:
+$\bm{P}^{t}=\pi_{\theta}(\bm{s}^{t}).$
+where the input $\bm{s}^{t}$ and output $\bm{P}^{t}$ are illustrated as follows.
 Input.
-As shown in Figure 1, we represent the state $\bm{s}^{t}\in\mathbb{R}^{|\mathcal{C}|\times(|\mathcal{G}|+2)}$ as a one-hot matrix, where each row of the matrix is the one-hot vector of the item created by corresponding creator connected with the corresponding trust degree.
-The length of the one-hot vector is $|\mathcal{G}|+2$, because there are cases where the creator dose not create item.
+As shown in Figure 1, we represent the state $\bm{s}^{t}\in\mathbb{R}^{|\mathcal{C}|\times(|\mathcal{G}|+1)}$ as a one-hot matrix, where each row of the matrix is the one-hot vector of the item created by corresponding creator.
+The length of the one-hot vector is $|\mathcal{G}|+1$, because there are cases where the creator dose not create item.
 Output.
@@
 Specifically, the output of actor is $\bm{P}=\pi_{\theta}(\bm{s}^{t})\in\mathbb{R}^{|\mathcal{C}|\times(\mathcal{G}|+1)}$, where each row of $\bm{P}$ is a distribution over the suggestion set for corresponding creator.
-Then, the suggestions sent to creators are sampled from the corresponding row of $\bm{P}$, which constitute the platform’s action $\bm{a}^{t}\in\mathbb{R}^{|\mathcal{C}|}$ as shown in Figure 1.
-After sending suggestions to creators, creators produce items and the genres of these new items with creators’ trust degree constitute the next state $\bm{s}^{t+1}$.
+The suggestions sent to creators are sampled from the corresponding row of $\bm{P}$, which constitute the platform’s action $\bm{a}^{t}\in\mathbb{R}^{|\mathcal{C}|}$.
+After sending suggestions to creators, creators produce items and the genres of these new items constitute the next state $\bm{s}^{t+1}$.
 After that, platform recommends items to each user and get the total users’ clicks $\mathcal{R}$ as reward.
-Finally, the data $b^{t}=(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ is then added to the reply buffer $\mathcal{B}$ which is used to train the actor and critic networks in the offline training phase.
-4.2.2. Offline Training
+Finally, the data $b^{t}=(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ is then added to the reply buffer $\mathcal{B}$ which is used to train the actor and critic networks in the training phase.
+4.2.2. Training Phase
 We use the collected reply buffer $\mathcal{B}$ to train the actor and critic networks.
 We denote the critic network by $V_{\phi}(\bm{s})$, which is parameterized by $\phi$.
-The critic network takes state $\bm{s}^{t}$ as input and outputs the estimated state value of $\bm{s}^{t}$, which is used to approximate the state value function $V^{\pi}(\bm{s})$ and help optimize the actor network.
+The critic network takes state $\bm{s}^{t}$ as input and outputs the estimated state value of $\bm{s}^{t}$, which is used to approximate the state value function $V^{\pi_{\theta}}(\bm{s})$ and help optimize the actor network.
 Actor update. To update the actor, we use the clipped surrogate objective (Schulman et al., 2017) as the loss function:
-(2) $L(\pi_{\theta})=-\sum_{b\in\mathcal{B}}\min\left(\pi^{\prime}(\bm{a}|\bm{s}),\text{clip}\left(\pi^{\prime}(\bm{a}|\bm{s}),1-\epsilon,1+\epsilon\right)\right)A^{\pi_{\theta}}(\bm{a},\bm{s}),$
-where $\pi^{\prime}(\bm{a}|\bm{s})=\frac{\pi_{\theta}(\bm{a}|\bm{s})}{\pi_{\theta_{old}}(\bm{a}|\bm{s})}$, $\text{clip}(x,l,r)=\max(\min(x,r),l)$ and $\epsilon$ is the clip hyperparameter.
+(2) $L(\pi_{\theta})=-\sum_{b\in\mathcal{B}}\min\left(\pi_{\theta}^{\prime}(\bm{a}|\bm{s}),\text{clip}\left(\pi_{\theta}^{\prime}(\bm{a}|\bm{s}),1-\epsilon,1+\epsilon\right)\right)A^{\pi_{\theta}}(\bm{a},\bm{s}),$
+where $\pi_{\theta}^{\prime}(\bm{a}|\bm{s})=\frac{\pi_{\theta}(\bm{a}|\bm{s})}{\pi_{\theta_{old}}(\bm{a}|\bm{s})}$, $\text{clip}(x,l,r)=\max(\min(x,r),l)$ and $\epsilon$ is the clip hyperparameter.
 $\pi_{\theta_{old}}$ is the actor network after the last update and remains unchanged when the current actor network $\pi_{\theta}$ is updated.
@@
 (3) $L^{TD}(V_{\phi})=\sum_{b\in\mathcal{B}}(\mathcal{R}_{t}+\gamma V_{\phi}(\bm{s}^{t+1})-V_{\phi}(\bm{s}^{t}))^{2}.$
-We determine whether our algorithm has converged based on the policy loss $L(\pi_{\theta})$.
-Specifically, we regard the algorithm as converged when the policy loss remains nearly flat, defined as a decrease of less than 1% over 10 consecutive episodes.
-4.2.3. Trust prediction
-We use maximum likelihood estimation (MLE) to predict creators’ trust degrees.
-The training dataset $\mathcal{D}=\{(i,y)\}$ is collected from the interactions between the platform and creators, where $i$ denotes the creator’s ID, and $y\in\{0,1\}$ is a binary label indicating whether creator $i$ follows the platform’s suggestion.
-A neural network parameterized by $\psi$, denoted as $P_{\psi}$, is employed to model the creators’ trust.
-The network takes a creator’s ID as input and outputs the estimated trust degree, i.e., $\hat{d}_{i}=P_{\psi}(i)$.
-The parameters $\psi$ are learned by minimizing the MLE loss:
-$L(P_{\psi})=-\sum_{(i,y)\in\mathcal{D}}\left[y\log(P_{\psi}(i))+(1-y)\log(1-P_{\psi}(i))\right].$
+We assess convergence based on the policy loss $L(\pi_{\theta})$.
+Specifically, the algorithm is considered converged when the policy loss remains approximately stable, operationalized as a decrease of less than 1% over 10 consecutive episodes.
 Figure 2. Comparison of user clicks in RS between our method and other baselines on two different datasets.
 5. Experiments
-In this section, we conduct extensive experiments to verify the effectiveness of our method in improving the long-term welfare of RS by addressing the following research questions:
-RQ1: Can our method improve the long-term welfare of RS?
+In this section, we conduct extensive experiments to evaluate the effectiveness of our method in improving the long-term welfare of RS.
+Specifically, we aim to address the following research questions::
+RQ1: Can our method improve the long-term welfare of RS through information revelation?
 RQ2: What is the inherent insight that our method can enchance the long-term welfare?
-RQ3: How robust our method is, and whether it can perform well under various settings?
-The source code of LoRe and datasets are available  [^3] https://anonymous.4open.science/r/LoRe-E792.
+RQ3: How robust is our method, and how does it perform under various settings?
+The source code of LoRe and the datasets used in our experiments are publicly available at https://anonymous.4open.science/r/LoRe-E5F0.
 5.1. Experimental Setups
@@
 Simulation setups.
-To better approximate real-world RS environments, we adopt the multi-stakeholder RS simulation platform proposed by Ye et al.
+To better simulate the real-world RS environments, we adopt the multi-stakeholder RS simulation platform proposed by Ye et al.
 (2025), which integrates LLM-based agents to simulate the dynamic behaviors of both users and creators.
 Recent studies have demonstrated that LLMs are capable of accurately mimicking user preferences and the decision-making of bounded-rational creators (Wang et al., 2024; Zhang et al., 2024; Wang et al., 2025; Ye et al., 2025).
-To test the robustness and generalization ability of our method under diverse environments, we also conduct additional experiments under environments where creators are simulated by non-LLM-based models (See Section 5.4).
-Besides, we consider the active behaviors of creators and users, as well as creator retention, following previous works (Mladenov et al., 2020; Xu et al., 2023; Ye et al., 2025): each creator $c_{i}\in\mathcal{C}$ and each user $u_{i}\in\mathcal{U}$ has probability $p_{c_{i}}$ and $p_{u_{i}}$ to be active, respectively.
+To evaluate the robustness and generalization ability of our method under diverse environments, we additionally conduct experiments under environments where creators are simulated using non-LLM-based models (See Section 5.4).
+Following previous works (Mladenov et al., 2020; Xu et al., 2023; Ye et al., 2025), we also consider the active behaviors of creators and users, as well as creator retention.
+Specifically, each creator $c_{i}\in\mathcal{C}$ and each user $u_{i}\in\mathcal{U}$ has probability $p_{c_{i}}$ and $p_{u_{i}}$ to be active, respectively.
 In each round, these probabilities determine whether a creator or user is active.
@@
 Moreover, a creator $c_{i}$ is assumed to leave the platform if no clicks are received in $l$ consecutive rounds of creation.
-Table 2. The content diversity (CD) and average number (AN) of active creators across different methods on different datasets.
-YouTube Amazon
-Method CD AN CD AN
-LoRe (ours) 3.259 16.250 3.462 14.710
-Reweight 3.223 15.625 2.744 10.530
-P-MMF 3.196 15.930 2.697 11.025
-FairRec 3.182 14.245 2.646 9.880
-Most History Click 3.197 14.730 2.805 11.610
-Most Click 3.186 13.840 2.847 11.775
-w/o signal 3.177 11.750 2.546 9.665
+Table 2.
+Comparisons of average user clicks, content diversity and active creators of our methods and baselines on YouTube and Amazon datasets.
+MHC represents the most history click and MC denotes the most click.
+We categorize baselines into three types: naive suggestion strategy, re-ranking methods, and reward adjust method.
+User Clicks Content Diversity Active Creators
+Methods YouTube Amazon YouTube Amazon YouTube Amazon
+w/o Signal $101.211\pm 3.941$ $29.55\pm 0.896$ $3.177\pm 0.006$ $2.546\pm 0.021$ $11.750\pm 0.153$ $9.665\pm 0.074$
+Naive MHC $105.831\pm 3.058$ $57.913\pm 1.828$ $3.197\pm 0.012$ $2.805\pm 0.006$ $4.730\pm 0.214 11$$1.610\pm 0.175$
+[…]
+Re-Ranking P-MMF $101.599\pm 1.263$ $47.794\pm 1.114$ $3.196\pm 0.009$ $2.697\pm 0.011$ $5.230\pm 0.120 11$$1.025\pm 0.097$
+FairRec $97.171\pm 3.396$ $38.860\pm 0.747$ $3.182\pm 0.011$ $2.646\pm 0.009$ $14.245\pm 0.153$ $9.880\pm 0.068$
+Reward Adjust Reweight $\underline{110.891}\pm 2.505$ $47.412\pm 1.290$ $\underline{3.223}\pm 0.004$ $2.744\pm 0.013$ $\underline{15.625}\pm 0.127$ $10.530\pm 0.082$
+Ours LoRe $\bm{127.113}\pm 2.589$ $\bm{81.36}\pm 0.981$ $\bm{3.259}\pm 0.012$ $\bm{3.462}\pm 0.007$ $\bm{16.250}\pm 0.148$ $\bm{14.710}\pm 0.164$
 Baselines.
-We compared our method with several baselines, including: w/o Signal: revealing no information to the creator at all, except for their own creation history; and some naive information revelation strategies: Most History Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres the creator has ever created; Most Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres.
-In addition, we compared our method with approaches that focus on recommendation algorithm design, such as P-MMF (Xu et al., 2023) and FairRec (Guo et al., 2023).
-We also included a baseline that influences creators by adjusting their rewards, namely Reweight (Yao et al., 2024).
+We compared our method with several baselines, including the following: w/o Signal: revealing no information to the creator at all, except for their own creation history.
+Some naive suggestion strategies are as follows.
+Most History Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres the creator has ever created; Most Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres.
+In addition to these naive suggestion strategies, we compared our method with approaches that focus on recommendation algorithm design, such as P-MMF (Xu et al., 2023) and FairRec (Guo et al., 2023).
+We also included a baseline that influences creators by adjusting their rewards rather than information revelation, namely Reweight (Yao et al., 2024).
 Metrics.
-Following previous works (Ye et al., 2025; Yao et al., 2024), we use the cumulative user click number to measure the long-term welfare.
-We adopt RecAgent (Wang et al., 2025) as the user simulator, where users can choose to click or skip when presented with an item.
-The number of clicks by all users in each round is recorded as the user welfare for that round.
-Figure 3. The number of genres created by each creator.
+Following previous works (Ye et al., 2025; Yao et al., 2024), we measure the long-term welfare of RS using the cumulative number of user clicks.
+We adopt RecAgent (Wang et al., 2025) as the user simulator, in which users decide whether to click on or skip a presented item.
+The total number of clicks across all users in each round is recorded as the user welfare for that round.
+Figure 3. Number of genres created per creator. The blue box denotes the mean number of genres created by each creator under LoRe, while the red box denotes the mean under the w/o Signal baseline.
 5.2. Effectiveness of LoRe (RQ1)
@@
 The platform employs DIN (Zhou et al., 2018) as the based recommendation model, and both user and creator agents are powered by LLaMA3-8B (Touvron et al., 2023).
-In addition, we initialize creators’ trust degrees using a truncated Gaussian distribution, reflecting real-world scenarios where agents exhibit moderate trust levels, and extremely high or low trust is uncommon (Teacy, 2006; Teacy et al., 2012).
-Denoted by $\mathcal{N}(\mu,\sigma;0,1)$ a truncated Gaussian distribution with mean $\mu$, variance $\sigma$, lower bound 0, and upper bound $1$.
-Then, each creator’s trust degree is randomly sampled from distribution $\mathcal{N}(0.5,1.0;0,1)$.
-Figure 2 shows the changes in cumulative user clicks per round as the number of rounds increases.
-LoRe consistently achieves higher user welfare than all baselines on both the YouTube and Amazon datasets.
-Compared with approaches that only adjust recommendation algorithms or creator rewards (e.g., PMMF, FairRec, Reweight), LoRe performs substantially better, indicating that influencing creator behavior through selective information revelation is more effective for improving long-term user welfare.
-Existing methods also show inconsistent performance across datasets.
-For instance, Reweight outperforms all other baselines on YouTube but performs worse than naive information revelation strategies on Amazon.
-This is primarily due to the more skewed user distribution in the Amazon dataset compared to the YouTube dataset, i.e., the majority of users in the Amazon dataset favor a particular genre of content, while fewer users favor other genres, which means that existing methods may significantly degrade in performance under such conditions.
-Naive information revelation strategies (Most History Click, Most Click) also display unstable performance: they outperform w/o Signal on Amazon but are similar to it on YouTube.
-This suggests that naive strategies lack robustness across recommendation environments.
-In contrast, LoRe consistently achieves the best long-term welfare, demonstrating that strategic information revelation is not only more effective than naive approaches but also more robust across different environments.
-Figure 4. Training time cost and the influence of trust degree.
+Figure 2 illustrates the evolution of user clicks per round as the number of interaction rounds increases, while Table 2 reports the average number of user clicks over $T$ rounds for all methods and datasets.
+Across both the YouTube and Amazon datasets, LoRe consistently achieves higher user welfare than all baselines.
+Compared with methods that focus solely on adjusting recommendation algorithms or creator rewards (e.g., PMMF, FairRec, and Reweight), LoRe exhibits substantially better performance, indicating that influencing creators’ creation behavior through information revelation is a more effective mechanism for improving long-term user welfare.
+Existing baselines also exhibit inconsistent performance across datasets.
+For example, Reweight outperforms other baselines on the YouTube dataset but performs worse than naive information revelation strategies on the Amazon dataset.
+This discrepancy is primarily attributable to the more skewed user preference distribution in the Amazon dataset, where a large proportion of users favor a single content genre, while relatively few users prefer other genres.
+Under such conditions, the performance of existing methods can degrade significantly.
+Naive information revelation strategies (Most History Click and Most Click) likewise demonstrate unstable performance: they outperform the w/o Signal baseline on the Amazon dataset but perform comparably on the YouTube dataset.
+This further indicates that naive strategies lack robustness across different recommendation environments.
+In contrast, LoRe consistently achieves the highest long-term welfare across all settings, demonstrating that carefully designed information revelation is not only more effective than naive approaches but also more robust across diverse environments.
 5.3. Inherent insight of LoRe (RQ2)
@@
 At the macro level, we examine the content diversity and the average number of active creators during the interactions among the platform, creators, and users.
-At the micro level, we investigate the variation in the number of content genres produced by individual creators under different methods.
+At the micro level, we investigate how the number of item genres produced by individual creators varies across different methods.
+Figure 4.
+Training time cost and the influence of trust degree.
+The left figure presents the number of convergence rounds and training time per rounds for LoRe under different agent number.
+The right figure shows the user clicks of LoRe under different trust degree.
 Figure 5. The performance of different methods under various creator simulation models.
@@
 This demonstrates that our approach enhances the diversity of RS compared with other baselines, which in turn indirectly improves user welfare.
-In our simulations, we assumed that a creator would exit after receiving no clicks for 10 consecutive rounds.
+In our simulations, a creator would exit after receiving no clicks for 10 consecutive rounds.
 We calculated the average number of active creators under different methods across datasets.
@@
 As shown in Figure 3, the x-axis represents the creator ID, and the y-axis indicates the number of distinct genres created by each individual.
-Obviously, compared to the baseline "w/o Signal", our approach leads to an increase in the number of genres per creator, which directly contributes to the overall improvement in content diversity.
+Obviously, compared to the baseline w/o Signal, our approach leads to an increase in the number of genres per creator, which directly contributes to the overall improvement in content diversity.
 5.4. Robustness Analysis (RQ3)
@@
 Even with 1,000 agents, the net training time per round is less than 4 seconds.
-These results confirm that our method scales effectively to large-scale, real-world scenarios.
+These results confirm that our method scales effectively to large-scale scenarios.
 Influence of creator simulation models.
-As discussed in Section 3.3, our method is capable of handling various creator simulation models.
-To verify this and evaluate its robustness, we conduct experiments across different creator simulation models.
-Specifically, we consider both LLM-based simulation models: CreAgent (Ye et al., 2025) (with Mistral-7B and Qwen2.5-7B as backbones), and heuristic-based models: Creator Feature Dynamics (CFD) (Lin et al., 2024; Zhan et al., 2021), where creators adjust their creation strategies using user feedback as gradients scaled by a learning rate, and SimuLine (Zhang et al., 2023), where creators’ subsequent creations are determined by probabilistic sampling based on the number of clicks in previous steps.
-Figure 5 presents the performance comparison between our method and the baselines under different creator simulation models.
-The results clearly show that, regardless of whether the simulation model is LLM-based or heuristic-based, our approach consistently outperforms all baselines.
+To evaluate our method’s robustness, we conduct experiments under multiple creator simulation models.
+Specifically, we consider both LLM-based simulation models: CreAgent (Ye et al., 2025) (with Mistral-7B and Qwen2.5-7B as backbones), and heuristic-based models: Creator Feature Dynamics (CFD) (Lin et al., 2024; Zhan et al., 2021), in which creators adjust their creation behavior using user feedback as gradients scaled by a learning rate, and SimuLine (Zhang et al., 2023), where creators’ creations are determined by probabilistic sampling based on the number of clicks in previous steps.
+When LLM-based simulation models are used, the platform’s suggestions can be directly incorporated into creators’ prompts, thereby influencing their behavior.
+However, this mechanism is not applicable to heuristic-based models.
+To simulate the effect of platform suggestions in such settings, inspired by  Prasad et al.
+(2023), we design a trust-based model.
+Specifically, each creator $c_{i}$ is assigned a trust degree $d_{i}\in[0,1]$, representing the probability that the creator follows the platform’s suggestion upon receiving it.
+In our experiments, creators’ trust degrees are initialized using a truncated Gaussian distribution to reflect real-world scenarios in which agents typically exhibit moderate trust, while extremely high or low trust levels are uncommon (Teacy, 2006; Teacy et al., 2012).
+Denoted by $\mathcal{N}(\mu,\sigma;0,1)$ a truncated Gaussian distribution with mean $\mu$, variance $\sigma$, lower bound 0, and upper bound $1$.
+Then, each creator’s trust degree is randomly sampled from distribution $\mathcal{N}(0.5,1.0;0,1)$.
+Figure 5 reports the performance comparison between our method and the baselines under different creator simulation models.
+The results show that, regardless of whether the simulation model is LLM-based or heuristic-based, our approach consistently outperforms all baselines.
 In contrast, existing recommendation-algorithm-based methods (i.e., P-MMF and FairRec) exhibit inconsistent performance across different creator simulation models.
@@
 This is because it is designed for a specific creator decision model and thus may fail to generalize well when the decision model changes.
-Table 3. The average user clicks achieved through combining LoRe with FairRec or P-MMF on YouTube and Amazon datasets.
-Method YouTube Amazon
-LoRe (ours) 137.4954 95.1499
-P-MMF 108.5688 51.2752
-FairRec 113.2789 46.1865
-LoRe + P-MMF 141.2263 101.9969
-LoRe + FairRec 146.0092 98.6177
 Influence of creators’ trust degree.
-Although platforms can provide various suggestions to creators, real-world creators differ in their degree of trust toward the platform.
-Highly trusting creators are more likely to follow the platform’s suggestions, whereas those with low trust are less likely to be influenced.
-In addition, in practice, creators’ trust levels may evolve over time according to the number of clicks they receive Prasad et al.
-(2023).
-To evaluate the effectiveness of our method under different levels of creator trust, we conduct experiments where creators’ trust degrees are sampled from distributions with varying expectations.
+When heuristic-based models are used to simulate creators’ behavior, we assume that each creator has a degree of trust on the platform.
+In real-world RS, creators naturally differ in their trust.
+Highly trusting creators are more likely to follow platform suggestions, whereas creators with low trust are less susceptible to such influence.
+Moreover, creators’ trust may evolve over time as a function of the feedback they receive(Prasad et al., 2023).
+To evaluate the effectiveness of our method under different levels of creator trust, we conduct experiments where creators’ initial trust degrees are sampled from distributions with different expectations.
 Specifically, we conduct experiments with $\mu\in\{0.4,0.6,0.8\}$ and $\sigma=1.0$.
@@
 Figure 4 (b) shows the user clicks under trust degree distributions with different expectations on the YouTube dataset.
-We compare our method with Reweight, which performs best among the baselines on YouTube.
+We compare our method with Reweight, since it performs best among the baselines on YouTube.
 The results show that although our method’s performance fluctuates slightly across different trust distributions, it consistently outperforms Reweight, demonstrating robustness to varying levels of creator trust.
 Furthermore, while performance of our method decreases slightly as the expected trust degree decreases, the decline is marginal.
-This trend is reasonable, as lower trust leads to fewer creators adopting the platform’s suggestions.
-These findings highlight the importance of platforms maintaining their credibility by providing beneficial suggestions to creators.
+Overall, these findings underscore the importance of maintaining platform credibility by providing consistently beneficial suggestions to creators.
+Figure 6. Comparison of user welfare with our method and baselines under the RS environment with two different recommendation models, i.e, MF and BPR, respectively.
+Influence of different recommendation algorithms.
+In real-world RS environments, platforms employ different recommendation algorithms to select content from the item pool for users.
+To verify the effectiveness of our method under various recommendation models, we conduct experiments using different recommendation algorithms as the platform’s recommendation policy.
+Specifically, we select two of the most widely used recommendation models: Matrix Factorization (MF) (Koren et al., 2009) and Bayesian Personalized Ranking (BPR) (Rendle et al., 2012).
+As illustrated in Figure 6, our method consistently achieves higher user welfare than the baselines across different recommendation algorithms.
+This result demonstrates that the effectiveness of LoRe is not tied to a specific recommendation model, but instead generalizes well across different algorithmic.
+Table 3. The average user clicks achieved through combining LoRe with FairRec or P-MMF on different datasets.
+Method YouTube Amazon
+P-MMF $108.569\pm 1.213$ $51.275\pm 0.984$
+FairRec $113.279\pm 1.146$ $46.187\pm 0.813$
+LoRe (ours) $137.495\pm 2.523$ $95.150\pm 0.687$
+LoRe + P-MMF $\underline{141.226}\pm 2.266$ $\bm{101.997}\pm 1.084$
+LoRe + FairRec $\bm{146.009}\pm 1.742$ $\underline{98.618}\pm 0.994$
 Incorporated with existing methods.
-We integrate our method with existing approaches that focus on recommendation algorithm design, such as P-MMF and FairRec.
-We exclude Reweight from this combination, as both LoRe and Reweight aim to enhance user welfare by influencing creators’ behavior.
-Table 3 presents the average user clicks achieved by combining our method with other approaches on the YouTube and Amazon datasets.
-It is evident that combining our method with these approaches further improves the performance.
-Moreover, LoRe + FairRec achieves better performance on the YouTube dataset, whereas LoRe + P-MMF performs better on the Amazon dataset.
-These results demonstrates that our approach is compatible with other methods and highlights its scalability.
+We integrate our method with existing approaches that focus on recommendation algorithm design, including P-MMF and FairRec.
+We exclude Reweight from this combination, as both LoRe and Reweight aim to enhance user welfare by influencing creators’ creation behavior.
+Table 3 presents the average number of user clicks achieved by combining our method with these approaches on the YouTube and Amazon datasets.
+The results show that integrating LoRe with these methods further improves performance across both datasets.
+In particular, LoRe + FairRec achieves higher user welfare on the YouTube dataset, whereas LoRe + P-MMF performs better on the Amazon dataset.
+These findings demonstrate that LoRe is compatible with existing recommendation algorithm and highlight its scalability.
+6. Social Impacts
+Our method effectively enhances long-term user welfare in recommendation systems.
+Beyond improving user satisfaction and engagement over time, it also contributes to the development of a healthier and more sustainable ecosystem for content creation and consumption.
+A potential concern is that platform might misuse such suggestion strategy to manipulate creators’ behavior, benefiting some creators at the expense of others.
+However, the proposed suggestion strategy does not compel creators’ actions, that is creators retain full autonomy over the content they create.
+Moreover, if the platform were to provide suggestions that systematically harm creators’ earnings, it would erode the platform’s credibility, reduce creators’ trust, and ultimately diminish user welfare.
+Thus, platforms lack incentives to issue harmful or manipulative suggestions.
 7. Conclusions
-In this paper, we reveal that influencing creators’ behaviors could be a better way to improve user welfare.
-Inspired by the idea of Bayesian persuasion, we model the platform as the sender and the creators as the receivers.
-To overcome the problem of bounded rationality assumption, we model the creators as bounded rational agents.
-Then we propose a novel method, LoRe, which aims to learn how to influence creators by revealing information to them.
-Lore introduces a reinforcement learning algorithm to address the limitations imposed by the strong assumption in Bayesian persuasion.
+In this paper, we demonstrate that influencing creators’ creation behavior could be a better way to improve long-term user welfare compared to methods relying on recommendation algorithm design.
+Inspired by the idea of cheap talk, we model the platform as the sender and creators as the receivers.
+To address the challenges of applying cheap talk to RS, we propose a novel method, LoRe, which formulates the dynamic interaction between the platform and creators as a MDP.
+LoRe employs a reinforcement learning algorithm to optimize the platform’s suggestion strategy.
 Extensive experiments on two real-world datasets demonstrate the effectiveness of our approach and show that it can be seamlessly integrated into the existing RS ecosystem, as well as combined with other methods that focus on recommendation algorithm design.
-Appendix
-Appendix A Examples of Content Suggestion in Real Applications
-Figure 6. Examples of creation suggestions in real-world applications across different short video platforms, each of which has hundreds of millions of users.
-As the number of creators on short-video platforms continues to grow, providing creation suggestions to guide creators, thereby indirectly improving user experience, has become a key component of platform operations.
-Figure 6 presents examples of creation suggestions on short-video content platforms, each of which typically serves hundreds of millions of users.
-Although current commercial platforms have begun to adopt creation suggestions, they are typically limited to simple suggestions, such as suggesting visual effects or the most trending topics.
-Effectively guiding creators in real-world industrial settings remains an open challenge.
-To bridge this gap, we aims to explore how to design more effective suggestion strategies in real-world platform environments to enhance overall user welfare.
-Appendix B Omitted Definitions in Section 4
-We respectively define the state value function $V^{\pi}(\bm{s}^{t})$, the state-action value function $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$as:
-$V^{\pi}(\bm{s}^{t})=\operatorname{\mathbb{E}}_{(\bm{a}^{t},\bm{s}^{t+1},\dots)\sim\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right],$
-$Q^{\pi}(\bm{s}^{t},\bm{a}^{t})=\operatorname{\mathbb{E}}_{(\bm{s}^{t+1},\bm{a}^{t+1},\dots)\sim\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right].$
-Appendix C More Detailed Experimental Setups
-Our experimental server is equipped with 28 Intel(R) Core(TM) i9-10940X CPUs, 128GB of RAM, and 4 NVIDIA RTX 4090 GPUs, providing a total of 96 GB of VRAM.
-Appendix D Additional Experiments
-D.1. Influence of Different Recommendation Models
-Figure 7. Comparison of user welfare with and without our method under the RS environment with two different recommendation models, i.e, MF and BPR, respectively.
-In real-world RS environments, platforms adopt different recommendation algorithms to select content from the item pool for users.
-To verify the effectiveness of our method under various recommendation models, we conduct experiments using different recommendation algorithms as the platform’s policy.
-We select two of the most widely used recommendation models: Matrix Factorization (MF) (Koren et al., 2009) and Bayesian Personalized Ranking (BPR) (Rendle et al., 2012).
-As illustrated in Figure 7, across different recommendation models, our method consistently achieves better user welfare compared to the w/o signal baseline.
-Specifically, when using MF as the recommendation model, the user welfare with information signaling is 1.18 times that of the no-signal case (at the last round of simulation).
-When using BPR, this number increases to 1.26 times.
-Appendix E Limitations
-In our work, we use large language models (LLMs) to simulate the behaviors of users and content creators.
-Although previous work has verified the effectiveness of the LLM-empowered simulation agents in simulating users click behavior and creator content creation behavior, there is limited research on whether such agents can effectively simulate creators’ responses to platform-provided suggestions.
-Therefore, how to simulate the behavior of real-world creators after receiving platform suggestions is also a key challenge and represents a promising direction for future research.
-Appendix F Social Impacts
-Our method can effectively enhance the long-term user welfare in RS.
-This not only improves user satisfaction and engagement over time but also contributes to the development of a healthier and more sustainable ecosystem for content consumption and creation.
-To the best of our knowledge, our method does not pose any adverse effects on society.